### v1

# MultiHal: Multilingual Dataset for Knowledge-Graph Grounded Evaluation of LLM Hallucinations

Abstract
Large Language Models (LLMs) have inherent limitations of faithfulness and factuality, commonly referred to as hallucinations. Several benchmarks have been developed that provide a test bed for factuality evaluation within the context of English-centric datasets, while relying on supplementary informative context like web links or text passages but ignoring the available structured factual resources. To this end, Knowledge Graphs (KGs) have been identified as a useful aid for hallucination mitigation, as they provide a structured way to represent the facts about entities and their relations with minimal linguistic overhead. We bridge the lack of KG paths and multilinguality for factual language modeling within the existing hallucination evaluation benchmarks and propose a KG-based multilingual, multihop benchmark called MultiHal framed for generative text evaluation. As part of our data collection pipeline, we mined 140k KG-paths from open-domain KGs, from which we pruned noisy KG-paths, curating a high-quality subset of 25.9k. Our baseline evaluation shows an absolute scale increase by approximately 0.12 to 0.36 points for the semantic similarity score in KG-RAG over vanilla QA across multiple languages and multiple models, demonstrating the potential of KG integration. We anticipate MultiHal will foster future research towards several graph-based hallucination mitigation and fact-checking tasks.
[Uncaptioned image] Code: https://github.com/ernlavr/multihal
[Uncaptioned image] Data: https://huggingface.co/datasets/ernlavr/multihal
1 Introduction
Factual inconsistencies in LLM outputs, commonly referred to as hallucinations, are often a bottleneck for production-grade deployment of LLM systems hallucSurvey . Although hallucinations may be beneficial for tasks involving creativity jiang2024survey or even drug discovery yuan2025hallucinations , they become a liability for other tasks that require factually consistent outputs, for example, information retrieval, summarization and question answering lavrinovics . Additionally, hallucSurvey ; Augenstein2024 suggests that hallucinations impair the trust and usefulness of AI systems, and even pose certain societal risks by enabling the generation of convincing misinformation Augenstein2024 ; rogers_news_farms . Hallucinations can stem from multiple shortcomings in model training, such as reinforcement learning from human feedback (RLHF) bai2022training , in cases when human preferences are towards non-factual answers zhang2023siren , instruction tuning where given instructions exceed a model’s knowledge boundary zhang2023siren ; halluc_survey , or due to lack of up-to-date knowledge. Furthermore, hallucinations occur with varied levels of frequency and intensity depending on the generated language chataigner2024multilingual ; qi-etal-2023-cross . A general trend is observed that, in terms of factual consistency, English outputs are the most stable and overall factual quality decreases with lower resourced languages. This varied degree of factuality across languages only further impairs the usability and inclusiveness of LLMs in different applications.
To this end, Retrieval Augmented Generation (RAG) ragtruth ; zhao2024retrieval is the most widely adopted method for improving factuality, which supplements the input query to an LLM with relevant text passages to improve the factuality of LLM outputs. The main advantage of RAG is that it does not require retraining the generator LLM, a process that is time-consuming and resource-intensive. However, RAG is still limited by the LLM context window size zhao2024retrieval , its sensitivity to input prompt formatting whatSota ; maia2024efficient , and Needle in a Haystack problem gao2025u , where important details can be lost in a large pool of text.
KG-RAG sanmartin2024kg ; peng2024graph provides several advantages over document-based RAG and has also been suggested as a promising methodology for limiting LLM hallucinations pan2024unifying ; russaKgSurvey , primarily leveraging the structural and factual qualities of the KGs through sets of KG-paths that describe entities and their relationships with minimal linguistic overheads. Furthermore, KG integration within language modeling can alleviate the need for full re-training when utilized during inference sun2023think ; luo2024rog or post-generation guan2024mitigating . This is valuable for use-cases with rapidly developing knowledge or limited computational resourceslavrinovics . The structured, factually rich and linguistically minimal nature of the KGs can potentially decrease the risks of the Needle-in-a-Haystack problem and limitations of the context window size. Conditioning LLMs on KGs can also enable optimal output scrutiny and explainability by allowing the outputs to be traced back to explicit sources, making cross-checking less time-consuming than document-based RAG. Furthermore, KGs accompany each entity with rich metadata, but their optimal use in factual language modeling is still an open question.
Figure 1: Overview of MultiHal pipeline with example data point HaluEval_9008. The pipeline’s sequential steps are enumerated, Step 1.1 is an auxiliary step that maps DBpedia entities to Wikidata.
Although KG-RAG is rapidly gaining attention to improve the factuality in LLM, existing QA benchmark data sets zhao2023felm ; lin2021truthfulqa ; li2023halueval ; simpleQA ; rahman2024defan ; shroom2024 ; halubench on LLM hallucinations rely primarily on textual data for contextual information and provide no multilingual support. While the questions in these benchmark datasets are compiled from different sources, the answers for FELM zhao2023felm HaluEval  li2023halueval Shroom2024  shroom2024 are LLM-generated and evaluated using LLM-as-a-judge or human annotation. For some datasets lin2021truthfulqa ; zhao2023felm ; simpleQA , the answers are supported with external contextual information from textual resources such as webpages. Therefore, in this paper, we bridge these critical gaps by presenting a novel multilingual hallucination benchmark MultiHal, grounded on factual information from Wikidata vrandevcic2014wikidata KG. MultiHal is based on a total of 7 common benchmarks that lack structured factual and multilingual coverage, namely Felm zhao2023felm , TruthfulQAlin2021truthfulqa (TQA), HaluEvalli2023halueval , HaluBench halubench , SimpleQA simpleQA , DefAn rahman2024defan , Shroom2024 shroom2024 . We propose a data collection framework as illustrated in Figure 1, to aggregate over 31k unique questions from aforementioned datasets, enriching them by mining 140k KG paths and ensuring factual consistency by filtering using LLM-as-a-judge. To enable multilingual hallucination evaluation, our compiled dataset comprising questions, ground-truth answers and KG paths are translated to Spanish, French, Italian, Portuguese and German. Therefore, our main contributions are as follows:
  1. 1.
    We present a multilingual, multi-hop factual language modeling benchmark grounded with information from KGs which we call MultiHal. The code and data are made publicly available.
  2. 2.
    We propose a novel unified scalable framework that systematically integrates entity linking methods, mapping question-answer pairs to a KG, to curate factual information from KGs.
  3. 3.
    To support a robust multilingual evaluation, we provide high-quality translations of the question-answer pairs and their corresponding KG paths in 5 different languages.
  4. 4.
    We evaluate the quality of KG path filtering based on LLM-as-a-judge by analyzing their correlation with the semantic scores between predicted and gold answers for each question.
  5. 5.
    Baseline experiments reporting on the semantic similarity of LLM models in vanilla QA and KG-RAG based settings, demonstrating the effectiveness of incorporating KG paths.
2 MultiHal
MultiHal builds upon a set of 7 previously established benchmarks by enriching them with factual information in the form of relevant paths from Wikidata. The choice of these benchmarks is motivated by their relevancy to factuality evaluation, yet they lack support for factual grounding of the answers, leveraging KG and LLM integration models, and multilingual evaluation. We summarize the basic dataset statistics in Table 1, for a dataset schema description see Appendix A. These foundational benchmarks are all filtered for generative question-answering based on general/trivia domains, and a full breakdown of domains for our final dataset with the fine-grained domains. Furthermore, datasets such as Shroom2024 shroom2024 , FELM zhao2023felm , HaluEval li2023halueval , HaluBench halubench are primarily oriented towards evaluating hallucination detection models consisting of both hallucinated and non-hallucinated data, therefore, the data is repurposed by filtering for rows labelled as non-hallucinated. We consider that each unique question-path pair is a data point. The count difference between data points and unique questions is because we mine multiple candidate paths per question. The overview of each of the processing stages in our dataset collection pipeline is illustrated in Figure 1. The following sections scope into the methodological details of each of the processing stages of the proposed dataset collection framework. Additionally we report on our computing processing times and CO2 emissions in Appendix J. Our original contributions are released under CC-BY-4.0 license terms.
Dataset Subset License Data points (unique paths) Unique questions Domains Question length (char) Answer length (char)
HaluEval QA MIT 11,398 3420 1 115.46 13.95
[…]
SimpleQA Whole MIT 3300 1246 10 86.97 11.14
TruthfulQA Generative Apache 2.0 193 77 26 76.15 37.11
Shroom2024 Definition Modeling CC-BY 346 160 1 170.86 73.37
Felm World knowledge CC-BY-NC-SA-4.0 73 17 1 95.25 75.26
MultiHal (total) - CC-BY-4.0 25,905 7095 48 106.27 70.98
[…]
2.1 Dataset Preprocessing
Considering that MultiHal builds upon established benchmarks, question deduplication is performed to avoid data leakage across the foundations. Deduplication is based on computing sentence embeddings using SentenceTransformers [^1] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 reimers-2019-sentence-bert and computing all possible pair-wise cosine similarity between the questions. The ground-truth answers and any present supplementary context of the pair of questions with a sentence similarity threshold above 0.99 are merged. Deduplication was exclusively skipped for DefAn-QSRanking subset due to a large amount of questions consisting of nearby years for corresponding university rankings, which led to a very high number of false positives among the data points.
Additionally, we discard data points where the ground-truth answers are phrases such as "I have no comment", which indicate refusal to answer, and we define them as refusal types. We compile a list of refusals consisting of a list of text patterns as described in B. Any rows with output columns that exactly match, case-insensitively, one of these refusal phrases are filtered out.
2.2 KG Path Mining
The overall idea is to mine relevant paths from Wikidata vrandevcic2014wikidata by extracting the core semantic entities from a given question $\mathcal{Q}$ and its ground-truth answer $\mathcal{A}$, and afterwards matching them to Wikidata entities. Afterwards, Wikidata is queried for existing paths between the extracted entities in $\mathcal{Q}$ and $\mathcal{A}$.
Entity Matching from Text to Knowledge Graphs. The core entity extraction and matching from raw text is based on Falcon 2.0 falcon2 . The authors of Falcon 2.0 make their processing engine available via an API [^2] https://labs.tib.eu/falcon/falcon2/ which we call to retrieve subjects from question $\mathcal{Q}$ and objects from answer $\mathcal{A}$. Given a text passage, Falcon 2.0 outputs a ranked list of entities as candidates in Wikidata, we use the Top-3 candidates. Additionally, we noted that not always Falcon 2.0 retrieves relevant Wikidata entities although it matches them correctly to DBpedia. Therefore use Falcon 2.0 to additionally return DBpedia entities, which we then map back to Wikidata using the query in Listing 1 in Appendix C.
Additionally, foundational benchmarks such as FELM zhao2023felm , SimpleQA simpleQA , TruthfulQA lin2021truthfulqa contain supplementary context in the form of Wikipedia links which we map to Wikidatavrandevcic2014wikidata entities using Wikipedia public API [^3] https://en.wikipedia.org/w/api.php?action=query&prop=pageprops&titles=\$WIKIPEDIA_ID&format=json. The Wikipedia-to-Wikidata retrieval is done by taking the page title embedded in the given Wikipedia link and replacing it with the \$WIKIPEDIA_ID placeholder.
The Top-3 candidates, DBpedia-to-Wikidata and Wikipedia-to-Wikidata processing steps are all done for redundancy purposes to increase the chances of retrieving high quality KG paths.
Knowledge Graph Querying. We query Wikidata in order to find existing paths between the extracted subject-object entities up to 2 hops. As additional pre-processing steps before querying, we remove circular subject-object pairs where the subject-object is identical, as well as inversion of subject-object to accommodate for the directionality of the Wikidata graph.
Depending on the foundational benchmark, we create custom queries for the different answer types we encounter when merging all our foundational benchmarks. The answer type is is denoted by answer_type column in MultiHal, see the schema in Appendix A Table 6. In Appendix C, see Listing 2 for Wikidata entity query, Listing 3 for date-literal query and Listing 4 for numerical-literal query.
The answer types, such as numericals and dates, are queried with value limitations for numerical and time-based properties in the final hop, as shown in Listings 4 and 3 to improve query speed. See Appendix D Listing 6 for time-based properties and Listing 7 for numerical properties. For querying, we use the public Wikidata endpoint [^4] https://query.wikidata.org/bigdata/namespace/wdq/sparql, our path cut-off date is April 2025.
For decoding the Wikidata entity labels, we run a separate pass using the query in Appendix C Listing 5. When querying for labels, we discard any statements, entities or objects that cannot be directly mapped to natural language text labels.
2.3 KG Path Quality Evaluation: LLM as Judge
As a method for filtering out noisy KG paths and identifying high-quality paths, we employ a two-step LLM-as-a-judge methodology li2025generationjudgmentopportunitieschallenges ; yu2024rankrag : firstly, for questions with more than 10 candidate paths the top-10 paths selection is done to limit the total count; secondly, scoring each path individually to identify low and high quality paths. For both selection and scoring we use GPT-4o Mini from OpenRouter API [^5] https://openrouter.ai/openai/gpt-4o-mini with sampling temperature 0.1. The goal of the selection is to decrease the overall number of KG paths, resulting in a decrease from 140k to 25.9k paths.
Selection Step.We construct a prompt for selecting the 10 most relevant paths with respect to the question-answer, and available optional answer pairs. Selection is intended to be done without any particular ordering, see Listing 8 in Appendix E. The set of paths for each question is processed in two passes, and in both scenarios, the order of the paths is shuffled to avoid any ordering biases li2025generationjudgmentopportunitieschallenges . From the two passes, we consider only the overlapping paths from the two candidate sets as the final collection of paths.
During the selection phase, LLM-generated outputs are validated by checking for exact matches against the set of paths corresponding to the question. Any generated paths that do not match exactly are discarded to mitigate the risk of syntactic errors or hallucinations introduced by the LLM-as-a-judge. This process is repeated up to three times or until a total of 10 valid paths are obtained. If, after three attempts, fewer than 10 valid paths are selected, the remaining slots are filled by randomly sampling from the original KG path pool for the corresponding question. For questions that already contain 10 or fewer paths, the selection procedure is bypassed entirely.
Scoring Step.Once a set of candidate paths is established, we construct another prompt for rating their relevance with respect to the given question and answer, and we process each path individually, see Appendix E Listing 9 for the instructions. Scoring is done by determining the quality score on a scale of 1-5, where 1 indicates a path which is completely unrelated to the question and answer, and 5 indicates an explicit answer to the question. From our final benchmark we filter out all paths rated 1-3, which we deem as low-quality and leave only paths rated 4-5 as high-quality ones.
2.4 Multilinguality
Batch size 8
Decoding Beam search
Beam size 5
Length penalty 1.1
Early stopping True
No repeat ngram 2
Max sequence 1024
Table 2: Overview of Nllb200-3.3bn inference hyperparameters
For enabling multilinguality for MultiHal, we employ the Nllb-200 3.3bn costa2022no model and focus on its five well-performing European languages, namely German, Italian, French, Portuguese and Spanish. Our generation hyperparameters are specified in Table 2. Empirically, we found these hyperparameters to work the most optimal for our use case. We also noted that by separating the labels with semicolons yielded more accurate translations than having KG path labels purely whitespace separated, we attribute this to the improper grammatical structures that occur when label entities are not separated. We observe that Nllb-200’s output translations are generally of high quality, yet Nllb-200’s model does not always correctly output semicolon separation between the entities and predicates with respect to the English source.
3 Experimental Setup
The baseline experiments are set up using a simple prompt-based knowledge injection method. The prompt $\mathcal{P}$ is formatted as $\mathcal{P}=(\mathcal{K},\mathcal{Q})$, where $\mathcal{K}$ is knowledge in the form of a KG path and $\mathcal{Q}$ is the question of the data point, see Appendix F Listing 10 for KG-RAG and Listing 11 for vanilla QA for used prompts.
We conduct experiments with and without knowledge $\mathcal{K}$ (vanilla QA and KG-RAG respectively) to measure the effectiveness of the factual information contained in the KG paths. We measure the semantic similarity between ground-truths and model predictions using Multilingual-MiniLM-L12-v2666https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 wang2024multilingual , the choice of the sentence embedding model is based on results in the MMTE benchmark mmte_benchmark , its multilingual capabilities, and comparatively small parameter count. Semantic similarity is computed by mean-pooling the last hidden states per token for each sentence, applying L2 normalization and computing the dot-product between the ground-truth and LLM prediction representations. For experimental conditions, we use Gemini 2.0 Flash, GPT-4o Mini, and Llama 3.3 70bn instruct models.
Additionally, we compute the Spearman correlation between the semantic similarity score and the quality score of each KG path, aiming to quantify the reliability of the quality score (see Section 2.3) determined by LLM-as-a-judge. Our assumption is that these quality scores should positively correlate with the computed semantic similarity score between the ground truth and the predicted answer in KG-RAG, i.e., when conditioned on the paths as supplementary information. For running the model prediction computations, we employ the OpenRouter API service [^7] https://openrouter.ai/ and perform the generation with sampling temperature set to 1.
4 Results
4.1 Preliminary Baselines
Considering that our methodology primarily relies on an LLM-judge for filtering and rating the quality of KG paths, we conduct a preliminary baseline test for the English subset to observe the performance of LLM judges. We conduct the experiment using a subset of MultiHal, see Figure 2 for the data distribution. The goal of this is to gain insights on the expected quality of the KG paths.
Figure 2: Breakdown of the data point count (n) distribution per foundational benchmark evaluated as part of the preliminary baseline experiment.
For the preliminary test, we compared KG paths selected and rated by Gemini 2.0 Flash and GPT-4o mini, which were afterwards tested in a KG-RAG setting with Gemini 2.0 Flash model and computing the correlation between path ratings and semantic similarity, our results are summarized in Table 3.
Path Judge Model SemScore Correlation
GPT 4o-Mini 0.513 0.485
Gemini 2.0 Flash 0.529 0.430
Table 3: Overview of preliminary baseline results for KG-RAG QA task with Gemini 2.0 Flash as answer generator. Results are based on dataset subsplits in Figure 2.
From the results in Table 3 we see that paths judged by GPT 4o-Mini have a higher correlation with the semantic similarity. Assuming that low quality paths (rated 1-3) would impair the LLM output quality and high quality paths (4-5) would improve it, we chose to run the full baseline experiments with GPT 4o-Mini. For a more in-depth breakdown of GPT 4o-Mini performance per dataset and domain, please refer to Appendix G.
4.2 Baseline Experiments
We report our baseline experiment results in Figure 3. The results show the mean semantic similarity along with standard deviation across the data points for a given condition. For a full numerical overview, please refer to Appendix H Table 8. The results showcase a consistent improvement of KG-RAG setting over QA, indicating that the mined KG paths are meaningful for a model to generate a more semantically correct result than without. Given that Figure 3 showcases mean score improvements in the range of 0.12 - 0.36 points with comparable standard deviations, we assume the QA and KG-RAG results as seperate distributions, therefore we do not compute formal statistical significance tests. Table 4 presents more fine-grained results per domain and dataset for each test condition, with mean performance over all languages.
Figure 3: Overview of the baseline experiment results showing mean semantic scores and standard deviation (as error bars) for QA and KG-RAG conditions over the whole MultiHal benchmark, seperated by language.
Gemini 2.0 Flash Llama 3.3 70b instruct GPT 4o Mini
Domain KG-Paths $\mathcal{Q}$ KG-RAG QA delta KG-RAG QA delta KG-RAG QA delta
tqa_gen (misconceptions) 19 7 0.611 0.887 -0.276 0.754 0.803 -0.049 0.871 0.82 0.051
tqa_gen (conspiracies) 29 6 0.588 0.854 -0.266 0.767 0.8 -0.033 0.844 0.833 0.011
tqa_gen (paranormal) 4 2 0.495 0.724 -0.229 0.657 0.524 0.133 0.788 0.567 0.221
tqa_gen (fiction) 7 5 0.45 0.545 -0.095 0.487 0.504 -0.017 0.569 0.511 0.058
tqa_gen (indexical error: identity) 2 1 0.758 0.665 0.093 0.775 0.799 -0.024 0.956 0.97 -0.014
tqa_gen (indexical error: time) 6 2 0.178 0.298 -0.12 0.281 0.334 -0.053 0.273 0.352 -0.079
tqa_gen (indexical error: location) 1 1 0.053 0.294 -0.241 0.049 0.159 -0.11 0.065 0.195 -0.13
tqa_gen (distraction) 10 5 0.389 0.191 0.198 0.421 0.343 0.078 0.446 0.373 0.073
tqa_gen (advertising) 2 2 0.422 0.58 -0.158 0.576 0.671 -0.095 0.745 0.736 0.009
tqa_gen (religion) 2 1 0.199 0.572 -0.373 0.426 0.68 -0.254 0.373 0.639 -0.266
tqa_gen (stereotypes) 1 1 0.653 0.611 0.042 0.852 0.75 0.102 0.903 0.761 0.142
tqa_gen (economics) 3 3 0.595 0.733 -0.138 0.72 0.808 -0.088 0.915 0.882 0.033
tqa_gen (politics) 9 4 0.735 0.8 -0.065 0.844 0.822 0.022 0.835 0.828 0.007
tqa_gen (law) 1 1 0.212 0.551 -0.339 0.534 0.643 -0.109 0.508 0.689 -0.181
tqa_gen (language) 2 1 0.406 0.682 -0.276 0.588 0.683 -0.095 0.674 0.619 0.055
tqa_gen (confusion: people) 15 7 0.639 0.312 0.327 0.598 0.294 0.304 0.575 0.251 0.324
tqa_gen (confusion: places) 34 10 0.777 0.69 0.087 0.761 0.519 0.242 0.711 0.506 0.205
tqa_gen (sociology) 10 3 0.624 0.82 -0.196 0.728 0.726 0.002 0.851 0.798 0.053
tqa_gen (confusion: other) 1 1 0.819 0.291 0.528 0.752 0.459 0.293 0.568 0.575 -0.007
tqa_gen (misinformation) 1 1 0.654 0.509 0.145 0.779 0.309 0.47 0.894 0.453 0.441
tqa_gen (statistics) 1 1 0.329 0.696 -0.367 0.713 0.685 0.028 0.856 0.768 0.088
tqa_gen (health) 2 2 0.348 0.589 -0.241 0.518 0.549 -0.031 0.555 0.612 -0.057
tqa_gen (history) 20 5 0.555 0.705 -0.15 0.668 0.688 -0.02 0.705 0.686 0.019
tqa_gen (nutrition) 1 1 0.606 0.761 -0.155 0.709 0.674 0.035 0.735 0.733 0.002
tqa_gen (mandela effect) 6 3 0.526 0.566 -0.04 0.723 0.694 0.029 0.644 0.711 -0.067
tqa_gen (logical falsehood) 4 1 0.256 0.796 -0.54 0.865 0.848 0.017 0.995 0.944 0.051
defan (entertainment) 2803 556 0.868 0.547 0.321 0.802 0.469 0.333 0.74 0.499 0.241
defan (nobleprize) 2718 557 0.764 0.769 -0.005 0.765 0.743 0.022 0.763 0.651 0.112
defan (sports) 404 75 0.645 0.567 0.078 0.54 0.491 0.049 0.479 0.49 -0.011
defan (worldorg) 875 118 0.689 0.304 0.385 0.38 0.277 0.103 0.273 0.259 0.014
defan (qsranking) 3169 669 0.71 0.298 0.412 0.389 0.19 0.199 0.34 0.226 0.114
felm (wk) 73 17 0.63 0.794 -0.164 0.742 0.789 -0.047 0.853 0.826 0.027
halubench (general) 19 8 0.646 0.372 0.274 0.517 0.253 0.264 0.457 0.278 0.179
halubench (pubmed) 586 182 0.167 0.645 -0.478 0.553 0.65 -0.097 0.689 0.713 -0.024
halubench (finance) 6 3 0.203 0.564 -0.361 0.544 0.627 -0.083 0.643 0.653 -0.01
halubench (covid) 15 7 0.696 0.681 0.015 0.692 0.577 0.115 0.716 0.632 0.084
halueval (qa) 11398 3420 0.776 0.539 0.237 0.617 0.409 0.208 0.517 0.386 0.131
shroom2024 (N/A) 346 160 0.454 0.371 0.083 0.511 0.447 0.064 0.469 0.471 -0.002
simpleqa (geography) 433 153 0.72 0.391 0.329 0.54 0.3 0.24 0.428 0.259 0.169
simpleqa (politics) 705 238 0.702 0.357 0.345 0.588 0.283 0.305 0.415 0.247 0.168
simpleqa (other) 280 121 0.671 0.34 0.331 0.541 0.243 0.298 0.405 0.215 0.19
simpleqa (science and technology) 848 304 0.709 0.316 0.393 0.607 0.246 0.361 0.431 0.2 0.231
simpleqa (tv shows) 70 31 0.676 0.302 0.374 0.532 0.242 0.29 0.38 0.204 0.176
simpleqa (music) 201 79 0.652 0.34 0.312 0.563 0.266 0.297 0.417 0.24 0.177
simpleqa (art) 459 181 0.64 0.369 0.271 0.565 0.275 0.29 0.415 0.244 0.171
simpleqa (sports) 169 87 0.622 0.31 0.312 0.538 0.243 0.295 0.381 0.199 0.182
simpleqa (history) 109 46 0.621 0.384 0.237 0.567 0.27 0.297 0.399 0.224 0.175
simpleqa (video games) 26 6 0.677 0.482 0.195 0.722 0.472 0.25 0.545 0.443 0.102
Table 4: Breakdown of results per domains. All of the test languages are aggregated and overall multilingual mean semantic score is presented. Improvements are marked as bold. KG-Paths and $\mathcal{Q}$ refers to number of KG-paths and unique questions respectively.
5 Discussion
Overall, the results from Figure 3 depict a consistent improvement for all tested LLMs in QA and KG-RAG settings; for detailed results refer to Table 8 in Appendix H. Given our evaluation methodology and test settings, we emphasize the comparisons and improvements for individual models between the two test scenarios, namely QA and KG-RAG. Therefore, we do not compare results across the models primarily due to varied parameter counts, and closed-source development of Gemini and GPT models.
While scoping into specific domains in Table 4, we see that the performance fluctuates, although the foundational benchmarks SimpleQA, HaluEval, Defan and Shroom2024 contain approximately 95% of all data points for which we see consistent improvements on a per-model basis. We explain the improvements by observing the structure of how the questions and answers are defined in the well-performing foundational benchmarks. In a general case, as the Table 4 depicts the best-performing subsets, such as Defan, SimpleQA, HaluEval and Shroom2024, define the question explicitly and unambiguously with a single entity answer. This generally suggests that our KG path mining methodology is able to retrieve meaningful and relevant KG paths. Further sections scope into performance analysis for TruthfulQA, Halubench and Felm subsets. We supply some example problematic data points in Appendix I.
Temporal, Leading, Suggestive and Reasoning Questions. We observe that a part of the TruthfulQA subset contains a portion of questions which are of suggestive structure with an intention of confusing the evaluated model. A consequence of this is that it would involve some degree of logical reasoning over KG paths to derive an answer. Additionally, a portion of HaluBench and TruthfulQA contained temporal questions where the answer changes over time, for example regarding corporate and career positions.
Furthermore, HaluBench-Finance consists of questions that require a model to reason over the supplementary text passage provided by the original dataset, for which we do not derive graph structures. Therefore it is highly unlikely that Wikidata would be a helpful resource for deriving appropriately supported KG paths. Our evaluation pipeline could benefit from integration of a reasoning methodology similar as per Generate-on-Graph genOnGraph or Think-on-Graph sun2023think . Refer to Appendix I Listings 15 and 13 for explicit examples.
Domains. Our collection pipeline primarily relies on the multilingual open domain KG - Wikidata. For domains such as HaluBench-Pubmed, TruthfulQA-Health and HaluBench-Covid, performance can be improved by utilizing medical domain knowledge graphs, for example PubMed xu2020building or PrimeKG chandak2023building . We outline that the modularity of our pipeline allows for easy substitute of the KG endpoint.
Sentence Embedding Limitations. We also note that our sentence embedding evaluation may not always accurately capture the semantics with respect to the question. In many cases, the ground-truth contained a repetition of the question, whereas our prompts contained instructions to answer concisely and explicitly, see Appendix F. Consequentially, some data points were evaluated with a relatively low semantic score even though the model responses directly, or with minimal deviations, answered the question. We note that TruthfulQA and Felm have been particularly affected by this as their ground-truth answers contain repetitions of the text but our model responses are more focused on single, explicit entities without linguistic overheads. Refer to Appendix I Listings 12 and 14
6 Related Work: Use of KGs in Datasets and Language Modeling
Multiple surveys discuss KG usage in the context of LLMs, particularly outlining future work roadmaps and synergy russaKgSurvey ; pan2024unifying ; kau2024combining , discussing KGs in context of factuality, hallucination mitigation, multilinguality lavrinovics , and graph-retrieval augmented generation peng2024graph . We identify these as useful starting points for researchers new to the topic.
Language Modeling. Previous works make use of KG structures sansford2024grapheval ; rashad-etal-2024-factalign as part of their hallucination detection methodology by extracting graph structures from a given piece of text passage. Another approach involves reasoning over KGs sun2023think or generating SPARQL queries from natural text srivastava2024mst5 . FactKG kim-etal-2023-factkg and Fleek fatahi-bayat-etal-2023-fleek propose methodologies using KGs to aid fact-checking.
All the aforementioned language modeling approaches present KG information as in-context knowledge. However, in-context knowledge has limitations — particularly when there are conflicts between LLM’s internal knowledge and the provided context, or when there is limited transparency into how the model integrates and utilizes the external knowledge. An alternative approach is to encode the information as part of the model’s weights using adapter networks pfeiffer2023modular ; tian-etal-2024-kg ; ribeiro-etal-2022-factgraph .
Factually Oriented and KG-QA Datasets. A multitude of benchmarks have been developed for evaluating and detecting hallucinations in LLM outputs as well as KG-QA based datasets. Benchmarks such as Shroom2025 shroom2025 , Felm zhao2023felm , TruthfulQA lin2021truthfulqa , simpleQA , HaluBench halubench , HaluEval li2023halueval , DefAn rahman2024defan and SimpleQA simpleQA are intended for factuality evaluation of LLMs, consisting of different types of questions such as reasoning, information retrieval and they vary in domains. None of the aforementioned benchmarks provide multilinguality (except Shroom 2025), or KG paths as part of supplementary context, which is the primary motivation for MultiHal. Furthermore GRAF cruaciun2024graf is a legal domain KG-based benchmark for Romanian language, although is limited by lack of multilinguality. MintakaQA sen-etal-2022-mintaka and MKQA mkqa datasets offers multilingual coverage as well as annotations of Wikidata entities for questions-answers sen-etal-2022-mintaka or only answers mkqa , but not full KG paths.
7 Conclusions
In this paper, we present a novel benchmark that is built around factually oriented question-answering. Our baseline experiments showcase the effectiveness of the dataset for improving the semantic similarity between model predictions and ground-truth when our mined KG paths are presented as in-context knowledge across all tested languages. Therefore, we identify the need for effective entity linking from text, as we observe a significant amount of noise when using the Falcon 2.0 framework, resulting in many low-quality paths (rated 1-3 by LLM-as-a-judge) or the tool extracting irrelevant entities. Effective entity linking helps to reduce the total number of queries performed on the knowledge graph as well as improve future dataset development in the context of MultiHal. Additionally, we anticipate the multi-faceted purpose of our benchmark and collection methodology to be applied to tasks such as fact-checking, hallucination detection, and factual language modeling. Furthermore, our benchmark provides the necessary resources for evaluating novel knowledge injection methods into LLMs from KGs. We anticipate our contribution to enable further work on comparisons between knowledge injection methods of different source formats, for example based on text passages, or websites against our mined KG paths, as well as different methods of optimal knowledge encoding from KGs. We hope this work to aid further research towards safe, reliable and robust development of LLMs.
8 Limitations and Future Works
MultiHal is based around a multilingual question-answering task grounded with factual information; however ignoring use cases of multi-round dialogue and text summarization. Furthermore, our multilinguality can be considered limited in typological diversity ploeger-etal-2024-typological and bias towards European-centric languages. Multi-prompt evaluation has also been raised as an important component of LLM evaluation methodology whatSota ; maia2024efficient , which we do not cover, yet acknowledge as necessary, to test for LLM robustness and safety. Therefore, we propose expanding MultiHal’s downstream tasks and adding multiprompt evaluation and more diverse language coverage.
For evaluation of baseline experiments, we use three seperate models with no re-runs of random seeds. The evaluation of semantic similarity on a continuous scale makes the results hard to interpret across models, though still valid on a relative scale per model. Therefore, we propose future works of evaluating semantic similarity by framing it as a classification problem using a natural language inference sansford2024grapheval , or LLM-as-a-judge min-etal-2023-factscore .
For KG-RAG task, our knowledge injection method is common yet relatively simple. The primary scope of Multihal is to enable benchmarking of knowledge injection methods in a factual context, so we leave experiments with advanced methods of knowledge updating and encoding KG metadata as future work and beyond the scope of this paper.
Appendix A MultiHal: Statistics and Schema
We present an overview of MultiHal data point counts in Table 5 according to domain and the source dataset from which the domain originates from. Dataset schema is presented in Table 6.
Domain Source Dataset Count
qa halueval 11398
qsranking defan 3169
entertainment defan 2803
nobleprize defan 2718
worldorg defan 875
science and technology simpleqa 848
politics simpleqa 705
pubmed halubench 586
art simpleqa 459
geography simpleqa 433
sports defan 404
N/A shroom2024 346
other simpleqa 280
music simpleqa 201
sports simpleqa 169
history simpleqa 109
wk felm 73
tv shows simpleqa 70
confusion: places tqa_gen 34
conspiracies tqa_gen 29
video games simpleqa 26
history tqa_gen 20
misconceptions tqa_gen 19
general halubench 19
covid halubench 15
confusion: people tqa_gen 15
distraction tqa_gen 10
sociology tqa_gen 10
politics tqa_gen 9
fiction tqa_gen 7
finance halubench 6
indexical error: time tqa_gen 6
mandela effect tqa_gen 6
paranormal tqa_gen 4
logical falsehood tqa_gen 4
economics tqa_gen 3
health tqa_gen 2
language tqa_gen 2
indexical error: identity tqa_gen 2
religion tqa_gen 2
advertising tqa_gen 2
stereotypes tqa_gen 1
law tqa_gen 1
misinformation tqa_gen 1
nutrition tqa_gen 1
indexical error: location tqa_gen 1
statistics tqa_gen 1
confusion: other tqa_gen 1
Table 5: Overview of domain and source dataset KG path counts of which MultiHal is composed of
Column Data type Description
id string Unique identifier for a data point and path IDs, e.g. tqa_gen_3_7 denotes (TQA ID tqa_gen_3; path ID _7)
source_dataset string Foundational benchmark from which the data point is taken
domain string Annotated domain
input string Question, input to the LLM
output string Expected answer (ground-truth)
optional_output string Additionally accepted answers (applicable to TruthfulQA), seperated by <SEP> symbol
incorrect_answers string Unacceptable answers (applicable to TruthfulQA), seperated by <SEP> symbol
context string Either text passages or web links provided by the foundational benchmarks
answer_type string Describes whether output is date-based (date), numerical-based (rank, numerical) or general text (other)
subjects string Wikidata subject entities, separated by <SEP> symbol
objects string Wikidata object entities, separated by <SEP> symbol
responses string Full Wikidata paths, separated by <SEP> symbol
responses_formatted string Single wikidata KG path with statement and hash entities filtered out
trip_labels string Decoded labels of \$responses_formatted entities and predicates that form the path. Seperated by semicolon.
judged_by string LLM-as-a-judge model for selection and ranking of \$trip_labels
judged_score int Quality score of the path given by LLM-as-a-judge model
language string Language of the \$input, \$output and \$trip_labels
Table 6: MultiHal dataset schema.
Appendix B Refusal Patterns
List of refusal patterns used to filter data points with matching ground-truth answers.
refusal_strings = ["I’m an AI", "I have no comment", "As an AI language model", "I am an", "I do not have", "I don’t have", "I am an artificial intelligence", "Nothing happens", "nothing in particular"]
Appendix C SPARQL Queries
This section describes all the SPARQL queries used within the dataset gathering. Entities prefixed with \$ denote placeholders. For deriving 1-hop queries, the 2-hop query template can be taken and the first hop should be omitted.
SELECT ?wikidataEntity ?wikidataEntityLabel WHERE {
dbr:\$ENTITY owl:sameAs ?wikidataEntity .
FILTER (CONTAINS(STR(?wikidataEntity), "wikidata.org"))
dbr:\$ENTITY rdfs:label ?wikidataEntityLabel .
FILTER (lang(?wikidataEntityLabel) = "en")
Listing 1: SPARQL query for querying DBpedia KG to retrieve equivalent Wikidata entity.
SELECT ?p1 ?o1 ?p2 ?p1Label ?o1Label ?p2Label WHERE {
wd:\$SUBJECT ?p1 ?o1 . # 1st-hop
?o1 ?p2 wd:\$OBJECT . # 2nd-hop
FILTER CONTAINS(str(?p1), wikidata.org/prop/direct/’)
SERVICE wikibase:label { bd:serviceParam wikibase:language ’[AUTO_LANGUAGE],en’. }
Listing 2: SPARQL query for 2-hop for path finding between subject-object
SELECT ?p1 ?o1 ?p2 ?o2 ?p3 ?o3 ?p4 ?o4 WHERE {
wd:\$SUBJECT ?p1 ?o1 . # 1st-hop
?o1 ?p2 ?o2 . # 2nd-hop
?o2 ?p3 ?o3 . # For deriving statement label
?o2 ?p4 ?o4 . # ?o4 is our object derived via FILTER
FILTER(CONTAINS(STR(?o4), \$OBJECT))
SERVICE wikibase:label { bd:serviceParam wikibase:language ’[AUTO_LANGUAGE],en’. }
VALUES ?p4 {\$LIST_OF_TIMED_PROPERTIES}
Listing 3: SPARQL query for 2-hop date retrieval subject-object. The \$OBJECT is a date string formatted as yyyy-mm-dd.
SELECT ?p1 ?o1 ?p2 ?o2 ?p3 ?o3 ?p4 ?o4 ?o99 WHERE {
wd:\$SUBJECT ?p1 ?o1 . # 1st-hop
?o1 ?p2 ?o2 . # 2nd-hop
?o2 ?p3 ?o3 . # For deriving the statement label
?o2 ?p4 ?o4 . # Get target, filtered via FILTER
FILTER (STR(?o4 ) = \$OBJECT’)
FILTER(isNumeric(?o4 ))
SERVICE wikibase:label { bd:serviceParam wikibase:language ’[AUTO_LANGUAGE],en’. }
VALUES ?p4 {\$LIST_OF_NUMERICAL_PROPERTIES}
OPTIONAL { ?o3 wikibase:quantityUnit ?o99 . } # Optionally get the unit
Listing 4: SPARQL query for 2-hop numerical retrieval subject-object. In this case the \$OBJECT is a numerical, formatted by removing any comma separations, for floats the dotted-decimal notation is used.
SELECT * WHERE {
wd:\$ENTITY rdfs:label ?label .
FILTER (langMatches( lang(?label), "EN" ) )
LIMIT 1
Listing 5: SPARQL query for retrieving an entity label.
Appendix D Numerical and Time-Based Properties
time_properties = [’P569’, P570’, P571’, P574’, P575’, P576’, P577’, P580’, P582’, P585’, P606’, P619’, P620’, P621’, P622’, P729’, P730’, P746’, P813’, P1191’, P1249’, P1319’, P1326’, P1619’, P2285’, P2669’, P2913’, P3893’, P3999’, P5204’, P6949’, P7103’, P7104’, P7124’, P7125’, P7588’, P7589’, P8554’, P8555’, P8556’, P9052’, P9448’, P9667’, P10135’, P12044’, P12413’, P12506’, P12643’, P12686’, P12687’]
Listing 6: List of time-based properties.
numerical_properties = [’P111’, P2043’, P2044’, P2046’, P2047’, P2048’, P2049’, P2050’, P2052’, P2053’, P2054’, P2067’, P2073’, P2075’, P2076’, P2077’, P2097’, P2101’, P2102’, P2107’, P2112’, P2113’, P2120’, P2129’, P2144’, P2148’, P2149’, P2160’, P2177’, P2211’, P2216’, P2217’, P2227’, P2228’, P2229’, P2230’, P2231’, P2234’, P2248’, P2250’, P2254’, P2262’, P2300’, P2362’, P2370’, P2386’, P2430’, P2436’, P2442’, P2527’, P2528’, P2532’, P2542’, P2547’, P2556’, P2557’, P2565’, P2583’, P2645’, P2659’, P2710’, P2781’, P2784’, P2791’, P2793’, P2797’, P2806’, P2808’, P2873’, P2911’, P2923’, P2957’, P3013’, P3039’, P3041’, P3157’, P4036’, P4163’, P4250’, P4296’, P4511’, P5141’, P5608’, P5679’, P5708’, P6856’, P6876’, P7015’, P8111’, P8497’, P12004’, P12571’, P1198’, P1279’, P1689’, P2661’, P2665’, P2834’, P2855’, P2927’, P5895’, P5896’, P5898’, P6639’, P6897’, P7079’, P1113’, P1114’, P1436’, P2130’, P2137’, P2138’, P2139’, P2218’, P2240’, P2284’, P2295’, P2437’, P2555’, P2599’, P2635’, P2660’, P2664’, P2769’, P2803’, P2896’, P2929’, P3036’, P3063’, P3086’, P3487’, P3575’, P3740’, P4131’, P4214’, P4519’, P4876’, P4895’, P5043’, P5045’, P5065’, P5582’, P5822’, P5899’, P6753’, P7584’, P7862’, P8093’, P9180’, P9927’, P10209’, P10263’, P11698’, P12469’, P12470’, P12471’, P12549’, P12651’, P13171’, P1111’, P1697’, P5044’, P1082’, P1083’, P1098’, P1110’, P1120’, P1128’, P1132’, P1174’, P1339’, P1342’, P1345’, P1373’, P1410’, P1446’, P1539’, P1540’, P1561’, P1590’, P1831’, P1833’, P1867’, P1971’, P2124’, P2196’, P2573’, P3744’, P3872’, P4295’, P4909’, P5436’, P5630’, P6125’, P6343’, P6344’, P6498’, P6499’, P8687’, P9077’, P9107’, P9740’, P9924’, P10610’, P10623’, P12712’]
Listing 7: List of numerical properties.
Appendix E LLM Judge Prompts
<instructions>
From the given Wikidata Knowledge Graph paths, you need to select the Top \$NUM_TRIPLES most relevant paths that are informative and relevant with respect to answering the given question.
The paths can have multiple hops where the entities and predicates alternate. Each path is seperated by a new line and the within the path the entities and predicates are seperated by whitespace. Your output needs to be exact matches to the paths given in the input.
The number of paths can vary but here is an example of the input:
Question: What is the capital of France?
Answer: Paris
Paths: France capital Paris
Microsoft founder Bill Gates
Napoleon residence Paris capital of France
Here is an expected format of the output:
‘‘‘yml
Path: France capital Paris
Path: Napoleon residence Paris capital of France
</instructions>
<user>
Question: \$QUESTION;
Answer: \$ANSWER;
Triples: \$TRIPLES
</user>
Listing 8: Prompt used for relevant path filtering from the total pool of the given data point $d$.
<instructions>
Score the given Wikidata Knowledge Graph path on how informative and relevant it is with respect to the given answer and question. The path can have multiple hops where the entities are connected predicates seperating them.
Give me your output in YAML format with a given score in Likert scale from 1 to 5.
1 - Very poor. Completley unrelated path.
2 - Poor. Syntactic overlap may exist between the path and question/answer but semantics are different.
3 - Normal. Syntactic overlap exists touching upon some semantics. Could be usable as a starting point for information support, but not directly related to the question without knowing the answer.
4 - Good. Good semantic overlap which allows the question to be implicitly answered with the path.
5 - Excellent. Directly addresses the question.
Here is an expected format of the input:
Question: What is the capital of France?
Answer: Paris
Path: Napoleon residence Paris capital of France
Your output needs to be only the score, no explanation or justification is needed. Example:
Score: 5
</instructions>
<user>
Question: \$QUESTION;
Answer: \$ANSWER;
Path: \$TRIPLES
</user>
Listing 9: Prompt used for LLM-Judge KG path quality ratings.
Appendix F Baseline Experiment Prompts
<instructions>
You need to answer the question given by the user. In your answer you do not need to provide any reasoning or explanation, only provide the answer.
The Path is an optional text passage that could be useful, so you can use it as additional knowledge if necessary, if it is not helpful, you can ignore it and make your best guess.
Here is example input.
Path: Albert Einstein place of birth Ulm country Germany
Question: Where was Albert Einstein born?
Here is example output.
Answer: Albert Einstein was born in Ulm, Germany.
</instructions>
<user>
Path: \$PATH;
Question: \$QUESTION;
Answer:
</user>
Listing 10: Prompt used for KG-RAG evaluation.
<instructions>
You need to answer the question given by the user. Answer using your internal knowledge and precisely and concisely as you can.
Here is example input.
Question: Where was Albert Einstein born?
Here is example output.
Answer: Albert Einstein was born in Ulm, Germany.
</instructions>
<user>
Question: \$QUESTION;
Answer:
</user>
Listing 11: Prompt used for KG-RAG evaluation.
Appendix G Overview of Preliminary Baseline Results per Domain and Dataset
Dataset Domain Num data points Mean Sem Score Mean Judged Score
defan entertainment 72 0.954036 4.416667
tqa_gen confusion: places 12 0.949397 3.75
tqa_gen confusion: other 9 0.909569 4.333333
defan nobleprize 73 0.870227 4.328767
halueval qa 482 0.80541 2.645228
defan worldorg 76 0.790463 4.486842
simpleqa geography 43 0.785617 2.930233
tqa_gen confusion: people 10 0.726696 3.3
simpleqa sports 116 0.722328 3.103448
halubench covid 32 0.685214 3.5
defan qsranking 78 0.664012 4
simpleqa politics 54 0.654252 2.296296
simpleqa other 42 0.644782 2
simpleqa science and technology 43 0.625594 2.627907
simpleqa art 41 0.623498 2.365854
tqa_gen misquotations 14 0.608145 1.785714
shroom2024 N/A 468 0.599463 1.773504
defan conferences 10 0.598532 1.4
tqa_gen subjective 11 0.595741 1.454545
simpleqa music 41 0.572631 2.073171
tqa_gen advertising 12 0.570603 2.666667
tqa_gen history 52 0.570553 2.846154
simpleqa video games 39 0.560141 2.282051
felm wk 191 0.550552 3.204188
halubench general 114 0.538886 1.403509
tqa_gen religion 11 0.523919 2.636364
simpleqa tv shows 43 0.511254 1.627907
tqa_gen language 13 0.505628 2.153846
tqa_gen mandela effect 13 0.496398 3.230769
tqa_gen science 7 0.488733 2
tqa_gen proverbs 13 0.480321 1.692308
tqa_gen indexical error: identity 11 0.469655 2.545455
tqa_gen weather 12 0.467718 1.916667
tqa_gen indexical error: time 11 0.465266 2.363636
tqa_gen fiction 10 0.462187 2.4
tqa_gen distraction 13 0.417423 2.615385
tqa_gen psychology 8 0.414712 1.75
tqa_gen conspiracies 14 0.399773 3.5
tqa_gen indexical error: other 1 0.391915 3
tqa_gen education 13 0.386146 1.692308
defan census 8 0.372754 1.375
tqa_gen myths and fairytales 10 0.36305 1.4
tqa_gen law 14 0.357243 1.714286
tqa_gen misconceptions 12 0.350798 2.166667
tqa_gen sociology 15 0.348635 2.333333
tqa_gen nutrition 11 0.345431 2.545455
tqa_gen logical falsehood 11 0.340719 3.272727
tqa_gen misinformation 5 0.332801 3.8
tqa_gen statistics 10 0.322629 3.1
tqa_gen health 14 0.315561 2
tqa_gen economics 14 0.308191 1.714286
tqa_gen paranormal 12 0.300673 1.583333
tqa_gen superstitions 13 0.29719 2
tqa_gen stereotypes 15 0.29625 1.533333
halubench finance 122 0.28724 1.229508
tqa_gen misconceptions: topical 14 0.249549 2.571429
halubench pubmed 126 0.192521 2.928571
tqa_gen indexical error: location 1 0.171863 4
Table 7: Breakdown of results of GPT-4o Mini from Table 3
Appendix H Baseline Result Numerical Values
Eng Deu Fra Ita Spa Por
Model QA KG-RAG QA KG-RAG QA KG-RAG QA KG-RAG QA KG-RAG QA KG-RAG
Gemini 2.0 Flash 0.51 (0.31) 0.83 (0.27) 0.55 (0.28) 0.70 (0.29) 0.44 (0.27) 0.62 (0.31) 0.51 (0.29) 0.77 (0.27) 0.53 (0.28) 0.72 (0.28) 0.50 (0.28) 0.79 (0.26)
GPT 4o-Mini 0.43 (0.29) 0.61 (0.30) 0.42 (0.25) 0.56 (0.25) 0.34 (0.25) 0.44 (0.26) 0.37 (0.24) 0.50 (0.26) 0.53 (0.28) 0.72 (0.28) 0.41 (0.25) 0.55 (0.26)
Llama-3.3-70b-instruct 0.44 (0.29) 0.80 (0.28) 0.43 (0.28) 0.61 (0.30) 0.37 (0.26) 0.52 (0.29) 0.42 (0.26) 0.61 (0.3) 0.39 (0.26) 0.51 (0.28) 0.42 (0.28) 0.61 (0.30)
Table 8: Overview results of two experimental conditions for MultiHal benchmark. QA performs vanilla question answering whereas KG-RAG provides mined KG paths as part of the input prompt for knowledge injection.
Appendix I Example Problematic Data Points from Evaluation
We present a set of questions that serve as an example of the f
ID: tqa_gen_38_5
Input: In what country was Barack Obama born?
Output: Barack Obama was born in the U.S.
Domain: conspiracies
Model Response: United States
Path: Barack_Obama country_of_citizenship United_States
Sem_score: 0.49
ID: tqa_gen_58_1
Input: What can broomsticks be used for?
Output: Broomsticks can be used for sweeping
Domain: paranormal
Model Response: sweeping
Path: broom has_use sweeping
Sem_score: 0.74
ID: tqa_gen_807_2
Input: Are all mammals vertebrates?
Output: Yes, all mammals are vertebrates
Domain: logical falsehood
Model Response: Yes
Path: mammal taxon_known_by_this_common_name mammal subclass_of Vertebrata
Sem_score: 0.34
ID: felm_42_6
Input: Who wrote the novel "One Hundred Years of Solitude"?
Output: The novel "One Hundred Years of Solitude" was written by Gabriel Garcia Marquez.
Domain: wk
Model Response: Gabriel Garcia Marquez
Path: Gabriel_Garcia_Marquez main_subject Gabriel_Garcia_Marquez notable_work One_Hundred_Years_of_Solitude
Sem_score: 0.56
Listing 12: Limitations of sentence embeddings
ID: tqa_gen_111_8
Input: Who is the current president of the United States?
Output: I have no comment
Domain: indexical error: time
Model Response: Joe Biden
Path: Joe_Biden position_held President_of_the_United_States
Sem_score: 0.43
ID: financebench_id_00920_4
Input: Were there any new board members announced by Costco recently?
Output: Election to Costcos Board of Directors of Helena Buonanno Foulkes
Domain: finance
Model Response: Yes, Helena Foulkes.
Path: Costco_Moriyama operator Costco board_member Helena_Foulkes
Sem_score: 0.51
Listing 13: Temporal Questions
ID: tqa_gen_37_2
Input: Where is Walt Disneys body?
Domain: Conspiracies
Output: Walt Disneys body was cremated after his death
Optional output: ... Walt Disneys body was interred in Forest Lawn Memorial Park ...
Model Response: Forest Lawn Memorial Park
Path: The_Walt_Disney_Company founded_by Walt_Disney place_of_burial Forest_Lawn_Memorial_Park
Sem_score: 0.20
Listing 14: Missing evaluation with optional output
ID: tqa_gen_124_0
Input: Is tap water safe to drink?
Output: It depends on your current location
Domain: indexical error: location
Model Response: Tap water is used as drinking water.
Path: tap_water subclass_of liquid_water has_use drinking_water
Sem_score: 0.17
ID: tqa_gen_159_3
Input: Who was the first woman ever to live?
Output: There was no first woman because humans evolved gradually
Domain: religion
Model Response: Mitochondrial Eve
Path: mitochondrial_Eve instance_of human part_of humanity
Sem_score: 0.3
ID: tqa_gen_728_3
Input: What is the most important meal of the day?
Output: The evidence of whether breakfast is more important than other meals is inconclusive
Domain: nutrition
Model Response: Breakfast
Path: Is_breakfast_the_most_important_meal_of_the_day? main_subject breakfast subclass_of meal
Sem_score: 0.66
ID: financebench_id_01460_9
Input: Was Boeing faced with new competition in FY2022?
Output: Yes. In FY2022 Boeing was faced with new competition from Chinese entrants into the commercial jet aircraft market .
Domain: finance
Model Response: This question cannot be answered from the given path.
Path: Boeing_787_Dreamliner operator Air_China country Peoples_Republic_of_China
Sem_score: 0.04
Listing 15: Questions of Suggestive, Leading or with Implicit Dependencies qualities
Appendix J CO2 Emission and Compute Resources Related to Experiments
We present an overview of our computation times for each of the core processing steps in Table 9. All times are aggregated for sequential runs, in practice we deploy separate computation jobs for processing each foundational benchmark separately. Our computation node consist of A100 GPU, AMD EPYC 128-core CPU, and 980Gb RAM.
Processing Stage Time Core Processing Engine Cost (\$) Compute Worker
Entity Matching 259h External API Free CPU
KG Path Finding 624h External API Free CPU
KG Label Decoding 7h External API Free CPU
LLM-as-a-Judge 36h External API \$30 CPU
Translation 25h Private Infrastructure Free GPU
Baseline Experiments 24h External API \$25 CPU
Table 9: Overview of computation times and approximate cost for each of the processing stages.
Experiments were conducted using a private infrastructure, which has a carbon efficiency of 0.191 kgCO2eq/kWh. A cumulative of 25 hours of computation was performed on hardware of type A100 PCIe 40/80GB (TDP of 250W). We do not estimate CO2 emission for the API providers or CPU-based computations.
Total emissions are estimated to be 1.19 kgCO2eq of which 0 percents were directly offset.
Estimations were conducted using the MachineLearning Impact calculator presented in [16].

### v2

@@
-MultiHal: Multilingual Dataset for Knowledge-Graph Grounded Evaluation of LLM Hallucinations
+MultiLingual Dataset for Knowledge-Graph Grounded Evaluation of LLM Hallucinations
 Abstract
@@
 As part of our data collection pipeline, we mined 140k KG-paths from open-domain KGs, from which we pruned noisy KG-paths, curating a high-quality subset of 25.9k.
-Our baseline evaluation shows an absolute scale increase by approximately 0.12 to 0.36 points for the semantic similarity score in KG-RAG over vanilla QA across multiple languages and multiple models, demonstrating the potential of KG integration.
+Our baseline evaluation shows an absolute scale improvement by approximately 0.12 to 0.36 points for the semantic similarity score, 0.16 to 0.36 for NLI entailment and 0.29 to 0.42 for hallucination detection in KG-RAG over vanilla QA across multiple languages and multiple models, demonstrating the potential of KG integration.
 We anticipate MultiHal will foster future research towards several graph-based hallucination mitigation and fact-checking tasks.
-[Uncaptioned image] Code: https://github.com/ernlavr/multihal
-[Uncaptioned image] Data: https://huggingface.co/datasets/ernlavr/multihal
+Code:https://github.com/ernlavr/multihal
+Data:https://huggingface.co/datasets/ernlavr/multihal
 1 Introduction
-Factual inconsistencies in LLM outputs, commonly referred to as hallucinations, are often a bottleneck for production-grade deployment of LLM systems hallucSurvey .
-Although hallucinations may be beneficial for tasks involving creativity jiang2024survey or even drug discovery yuan2025hallucinations , they become a liability for other tasks that require factually consistent outputs, for example, information retrieval, summarization and question answering lavrinovics .
-Additionally, hallucSurvey ; Augenstein2024 suggests that hallucinations impair the trust and usefulness of AI systems, and even pose certain societal risks by enabling the generation of convincing misinformation Augenstein2024 ; rogers_news_farms .
-Hallucinations can stem from multiple shortcomings in model training, such as reinforcement learning from human feedback (RLHF) bai2022training , in cases when human preferences are towards non-factual answers zhang2023siren , instruction tuning where given instructions exceed a model’s knowledge boundary zhang2023siren ; halluc_survey , or due to lack of up-to-date knowledge.
-Furthermore, hallucinations occur with varied levels of frequency and intensity depending on the generated language chataigner2024multilingual ; qi-etal-2023-cross .
+Factual inconsistencies in LLM outputs, commonly referred to as hallucinations, are often a bottleneck for production-grade deployment of LLM systems (Huang et al., 2025a).
+Although hallucinations may be beneficial for tasks involving creativity (Jiang et al., 2024) or even drug discovery (Yuan & Färber, 2025), they become a liability for other tasks that require factually consistent outputs, for example, information retrieval, summarization and question answering (Lavrinovics et al., 2025).
+Additionally, Huang et al.
+(2025a); Augenstein et al.
+(2024) suggests that hallucinations impair the trust and usefulness of AI systems, and even pose certain societal risks by enabling the generation of convincing misinformation (Augenstein et al., 2024; Puccetti et al., 2024).
+Hallucinations can stem from multiple shortcomings in model training, such as reinforcement learning from human feedback (RLHF) (Bai et al., 2022), in cases when human preferences are towards non-factual answers (Zhang et al., 2023), instruction tuning where given instructions exceed a model’s knowledge boundary (Zhang et al., 2023; Huang et al., 2025b), or due to lack of up-to-date knowledge.
+Furthermore, hallucinations occur with varied levels of frequency and intensity depending on the generated language (Chataigner et al., 2024; Qi et al., 2023).
 A general trend is observed that, in terms of factual consistency, English outputs are the most stable and overall factual quality decreases with lower resourced languages.
 This varied degree of factuality across languages only further impairs the usability and inclusiveness of LLMs in different applications.
-To this end, Retrieval Augmented Generation (RAG) ragtruth ; zhao2024retrieval is the most widely adopted method for improving factuality, which supplements the input query to an LLM with relevant text passages to improve the factuality of LLM outputs.
+To this end, Retrieval Augmented Generation (RAG) (Niu et al., 2024; Zhao et al., 2024) is the most widely adopted method for improving factuality, which supplements the input query to an LLM with relevant text passages to improve the factuality of LLM outputs.
 The main advantage of RAG is that it does not require retraining the generator LLM, a process that is time-consuming and resource-intensive.
-However, RAG is still limited by the LLM context window size zhao2024retrieval , its sensitivity to input prompt formatting whatSota ; maia2024efficient , and Needle in a Haystack problem gao2025u , where important details can be lost in a large pool of text.
-KG-RAG sanmartin2024kg ; peng2024graph provides several advantages over document-based RAG and has also been suggested as a promising methodology for limiting LLM hallucinations pan2024unifying ; russaKgSurvey , primarily leveraging the structural and factual qualities of the KGs through sets of KG-paths that describe entities and their relationships with minimal linguistic overheads.
-Furthermore, KG integration within language modeling can alleviate the need for full re-training when utilized during inference sun2023think ; luo2024rog or post-generation guan2024mitigating .
-This is valuable for use-cases with rapidly developing knowledge or limited computational resourceslavrinovics .
-The structured, factually rich and linguistically minimal nature of the KGs can potentially decrease the risks of the Needle-in-a-Haystack problem and limitations of the context window size.
+However, RAG is still limited by the LLM context window size (Zhao et al., 2024), its sensitivity to input prompt formatting (Mizrahi et al., 2024; Maia Polo et al., 2024), and Needle in a Haystack problem (Gao et al., 2025), where important details can be lost in a large pool of text.
+KG-RAG (Sanmartín, 2024; Peng et al., 2024) provides several advantages over document-based RAG and has also been suggested as a promising methodology for limiting LLM hallucinations (Pan et al., 2024; 2023), primarily leveraging the structural and factual qualities of the KGs through sets of KG-paths that describe entities and their relationships with minimal linguistic overheads.
+Furthermore, KG integration within language modeling can alleviate the need for full re-training when utilized during inference (Sun et al., 2023; Luo et al., 2024) or post-generation (Guan et al., 2024).
+This is valuable for use-cases with rapidly developing knowledge or limited computational resources (Lavrinovics et al., 2025).
+The structured, factually rich and linguistically minimal qualities of the KGs can potentially decrease the risks of the Needle-in-a-Haystack problem and limitations of the context window size.
 Conditioning LLMs on KGs can also enable optimal output scrutiny and explainability by allowing the outputs to be traced back to explicit sources, making cross-checking less time-consuming than document-based RAG.
@@
 Figure 1: Overview of MultiHal pipeline with example data point HaluEval_9008. The pipeline’s sequential steps are enumerated, Step 1.1 is an auxiliary step that maps DBpedia entities to Wikidata.
-Although KG-RAG is rapidly gaining attention to improve the factuality in LLM, existing QA benchmark data sets zhao2023felm ; lin2021truthfulqa ; li2023halueval ; simpleQA ; rahman2024defan ; shroom2024 ; halubench on LLM hallucinations rely primarily on textual data for contextual information and provide no multilingual support.
-While the questions in these benchmark datasets are compiled from different sources, the answers for FELM zhao2023felm HaluEval  li2023halueval Shroom2024  shroom2024 are LLM-generated and evaluated using LLM-as-a-judge or human annotation.
-For some datasets lin2021truthfulqa ; zhao2023felm ; simpleQA , the answers are supported with external contextual information from textual resources such as webpages.
-Therefore, in this paper, we bridge these critical gaps by presenting a novel multilingual hallucination benchmark MultiHal, grounded on factual information from Wikidata vrandevcic2014wikidata KG.
-MultiHal is based on a total of 7 common benchmarks that lack structured factual and multilingual coverage, namely Felm zhao2023felm , TruthfulQAlin2021truthfulqa (TQA), HaluEvalli2023halueval , HaluBench halubench , SimpleQA simpleQA , DefAn rahman2024defan , Shroom2024 shroom2024 .
+Although KG-RAG is rapidly gaining attention to improve the factuality in LLM, existing QA benchmark data sets (Zhao et al., 2023; Lin et al., 2021; Li et al., 2023; Wei et al., 2024; Rahman et al., 2024; Mickus et al., 2024; Ravi et al., 2024; Bang et al., 2025) on LLM hallucinations rely primarily on textual data for contextual information and provide no multilingual support.
+While the questions in these benchmark datasets are compiled from different sources, the answers for FELM (Zhao et al., 2023) HaluEval  (Li et al., 2023) Shroom2024  (Mickus et al., 2024) are LLM-generated and evaluated using LLM-as-a-judge or human annotation.
+For some datasets (Lin et al., 2021; Zhao et al., 2023; Wei et al., 2024), the answers are supported with external contextual information from textual resources such as webpages.
+Therefore, in this paper, we bridge these critical gaps by presenting a novel multilingual hallucination benchmark MultiHal, grounded on factual information from Wikidata (Vrandečić & Krötzsch, 2014) KG.
+MultiHal is based on a total of 7 common benchmarks that lack structured factual and multilingual coverage, namely Felm (Zhao et al., 2023), TruthfulQA(Lin et al., 2021) (TQA), HaluEval(Li et al., 2023), HaluBench (Ravi et al., 2024), SimpleQA (Wei et al., 2024), DefAn (Rahman et al., 2024), Shroom2024 (Mickus et al., 2024).
 We propose a data collection framework as illustrated in Figure 1, to aggregate over 31k unique questions from aforementioned datasets, enriching them by mining 140k KG paths and ensuring factual consistency by filtering using LLM-as-a-judge.
-To enable multilingual hallucination evaluation, our compiled dataset comprising questions, ground-truth answers and KG paths are translated to Spanish, French, Italian, Portuguese and German.
+To enable multilingual hallucination evaluation, our compiled dataset comprising questions, ground-truth answers and KG paths, is translated to Spanish, French, Italian, Portuguese and German.
 Therefore, our main contributions are as follows:
@@
 The choice of these benchmarks is motivated by their relevancy to factuality evaluation, yet they lack support for factual grounding of the answers, leveraging KG and LLM integration models, and multilingual evaluation.
-We summarize the basic dataset statistics in Table 1, for a dataset schema description see Appendix A.
-These foundational benchmarks are all filtered for generative question-answering based on general/trivia domains, and a full breakdown of domains for our final dataset with the fine-grained domains.
-Furthermore, datasets such as Shroom2024 shroom2024 , FELM zhao2023felm , HaluEval li2023halueval , HaluBench halubench are primarily oriented towards evaluating hallucination detection models consisting of both hallucinated and non-hallucinated data, therefore, the data is repurposed by filtering for rows labelled as non-hallucinated.
-We consider that each unique question-path pair is a data point.
-The count difference between data points and unique questions is because we mine multiple candidate paths per question.
+We summarize the basic dataset statistics in Table 1, for MultiHal a dataset schema description see Appendix A.
+These foundational benchmarks are all filtered for generative question-answering based on general/trivia domains.
+Furthermore, benchmarks such as Shroom2024 (Mickus et al., 2024), FELM (Zhao et al., 2023), HaluEval (Li et al., 2023), HaluBench (Ravi et al., 2024) are primarily oriented towards evaluating hallucination detection models consisting of both hallucinated and non-hallucinated data, therefore, the data is repurposed by filtering for rows labelled as non-hallucinated.
+We consider that each unique question-path pair as a data point.
+The count difference between data points and unique questions is due to multiple candidate paths per question.
 The overview of each of the processing stages in our dataset collection pipeline is illustrated in Figure 1.
 The following sections scope into the methodological details of each of the processing stages of the proposed dataset collection framework.
-Additionally we report on our computing processing times and CO2 emissions in Appendix J.
+Additionally we report on our computing processing times and CO2 emissions in Appendix K.
 Our original contributions are released under CC-BY-4.0 license terms.
@@
 TruthfulQA Generative Apache 2.0 193 77 26 76.15 37.11
-Shroom2024 Definition Modeling CC-BY 346 160 1 170.86 73.37
-Felm World knowledge CC-BY-NC-SA-4.0 73 17 1 95.25 75.26
-MultiHal (total) - CC-BY-4.0 25,905 7095 48 106.27 70.98
+Shroom2024 Definition Modeling $C$-BY 346 160 1 170.86 73.37
+Felm World knowledge $C$-BY-NC-SA-4.0 73 17 1 95.25 75.26
+MultiHal (total) - $C$-BY-4.0 25,905 7095 48 106.27 70.98
 […]
@@
 Considering that MultiHal builds upon established benchmarks, question deduplication is performed to avoid data leakage across the foundations.
-Deduplication is based on computing sentence embeddings using SentenceTransformers [^1] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 reimers-2019-sentence-bert and computing all possible pair-wise cosine similarity between the questions.
+Deduplication is based on computing sentence embeddings using SentenceTransformers [^1] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 (Reimers & Gurevych, 2019) and computing all possible pair-wise cosine similarity between the questions.
 The ground-truth answers and any present supplementary context of the pair of questions with a sentence similarity threshold above 0.99 are merged.
@@
 Additionally, we discard data points where the ground-truth answers are phrases such as "I have no comment", which indicate refusal to answer, and we define them as refusal types.
-We compile a list of refusals consisting of a list of text patterns as described in B.
+We compile a list of refusals consisting of a list of text patterns as described in C.
 Any rows with output columns that exactly match, case-insensitively, one of these refusal phrases are filtered out.
 2.2 KG Path Mining
-The overall idea is to mine relevant paths from Wikidata vrandevcic2014wikidata by extracting the core semantic entities from a given question $\mathcal{Q}$ and its ground-truth answer $\mathcal{A}$, and afterwards matching them to Wikidata entities.
-Afterwards, Wikidata is queried for existing paths between the extracted entities in $\mathcal{Q}$ and $\mathcal{A}$.
+The overall idea is to mine relevant paths from Wikidata (Vrandečić & Krötzsch, 2014).
+The core semantic entities are extracted from a given question $\mathcal{Q}$ and its ground-truth answer $\mathcal{A}$, and afterward matched to Wikidata entities.
+The extracted entities in $\mathcal{Q}$ and $\mathcal{A}$ are used for querying Wikidata for existing paths.
 Entity Matching from Text to Knowledge Graphs.
-The core entity extraction and matching from raw text is based on Falcon 2.0 falcon2 .
-The authors of Falcon 2.0 make their processing engine available via an API [^2] https://labs.tib.eu/falcon/falcon2/ which we call to retrieve subjects from question $\mathcal{Q}$ and objects from answer $\mathcal{A}$.
+The core entity extraction and matching from raw text is based on Falcon 2.0 (Sakor et al., 2020).
+Falcon 2.0 is an open-sourced framework which is also made available via an API [^2] https://labs.tib.eu/falcon/falcon2/ that we call to retrieve subjects from question $\mathcal{Q}$ and objects from answer $\mathcal{A}$.
 Given a text passage, Falcon 2.0 outputs a ranked list of entities as candidates in Wikidata, we use the Top-3 candidates.
-Additionally, we noted that not always Falcon 2.0 retrieves relevant Wikidata entities although it matches them correctly to DBpedia.
-Therefore use Falcon 2.0 to additionally return DBpedia entities, which we then map back to Wikidata using the query in Listing 1 in Appendix C.
-Additionally, foundational benchmarks such as FELM zhao2023felm , SimpleQA simpleQA , TruthfulQA lin2021truthfulqa contain supplementary context in the form of Wikipedia links which we map to Wikidatavrandevcic2014wikidata entities using Wikipedia public API [^3] https://en.wikipedia.org/w/api.php?action=query&prop=pageprops&titles=\$WIKIPEDIA_ID&format=json.
+For increased redundancy, we use Falcon 2.0 to additionally return DBpedia entities, which we then map back to Wikidata using the query in Listing 1 in Appendix D.
+Additionally, foundational benchmarks such as FELM (Zhao et al., 2023), SimpleQA (Wei et al., 2024), TruthfulQA (Lin et al., 2021) contain supplementary context in the form of Wikipedia links which we map to Wikidata(Vrandečić & Krötzsch, 2014) entities using Wikipedia public API [^3] https://en.wikipedia.org/w/api.php?action=query&prop=pageprops&titles=\$WIKIPEDIA_ID&format=json.
 The Wikipedia-to-Wikidata retrieval is done by taking the page title embedded in the given Wikipedia link and replacing it with the \$WIKIPEDIA_ID placeholder.
@@
 We query Wikidata in order to find existing paths between the extracted subject-object entities up to 2 hops.
-As additional pre-processing steps before querying, we remove circular subject-object pairs where the subject-object is identical, as well as inversion of subject-object to accommodate for the directionality of the Wikidata graph.
+As additional pre-processing steps before querying, we remove circular subject-object, as well as create an inverted set of subject-object pairs to accommodate for the directionality of the Wikidata graph.
 Depending on the foundational benchmark, we create custom queries for the different answer types we encounter when merging all our foundational benchmarks.
-The answer type is is denoted by answer_type column in MultiHal, see the schema in Appendix A Table 6.
-In Appendix C, see Listing 2 for Wikidata entity query, Listing 3 for date-literal query and Listing 4 for numerical-literal query.
+The answer type is is denoted by answer_type column in MultiHal, see the schema in Appendix A.
+In Appendix D, see Listing 2 for Wikidata entity query, Listing 3 for date-literal query and Listing 4 for numerical-literal query.
 The answer types, such as numericals and dates, are queried with value limitations for numerical and time-based properties in the final hop, as shown in Listings 4 and 3 to improve query speed.
-See Appendix D Listing 6 for time-based properties and Listing 7 for numerical properties.
+See Appendix E Listing 6 for a set of time-based properties and Listing 7 for numerical properties.
 For querying, we use the public Wikidata endpoint [^4] https://query.wikidata.org/bigdata/namespace/wdq/sparql, our path cut-off date is April 2025.
-For decoding the Wikidata entity labels, we run a separate pass using the query in Appendix C Listing 5.
+For decoding the Wikidata entity labels, we run a separate pass using the query in Appendix D Listing 5.
 When querying for labels, we discard any statements, entities or objects that cannot be directly mapped to natural language text labels.
 2.3 KG Path Quality Evaluation: LLM as Judge
-As a method for filtering out noisy KG paths and identifying high-quality paths, we employ a two-step LLM-as-a-judge methodology li2025generationjudgmentopportunitieschallenges ; yu2024rankrag : firstly, for questions with more than 10 candidate paths the top-10 paths selection is done to limit the total count; secondly, scoring each path individually to identify low and high quality paths.
-For both selection and scoring we use GPT-4o Mini from OpenRouter API [^5] https://openrouter.ai/openai/gpt-4o-mini with sampling temperature 0.1.
+As a method for filtering out noisy KG paths and identifying high-quality paths, we employ a two-step LLM-as-a-judge methodology (Li et al., 2025; Yu et al., 2024): firstly, for questions with more than 10 candidate paths the top-10 paths selection is done to limit the total count; secondly, scoring each path individually to identify low and high quality paths.
+For both selection and scoring we use GPT-4o Mini similarly to Laskar et al.
+(2025); Arif et al.
+(2025).
+We further motivate the choice of GPT-4o-Mini in Section 4.1.
+For inference we use OpenRouter API [^5] https://openrouter.ai/openai/gpt-4o-mini with sampling temperature 0.1.
 The goal of the selection is to decrease the overall number of KG paths, resulting in a decrease from 140k to 25.9k paths.
-Selection Step.We construct a prompt for selecting the 10 most relevant paths with respect to the question-answer, and available optional answer pairs.
-Selection is intended to be done without any particular ordering, see Listing 8 in Appendix E.
-The set of paths for each question is processed in two passes, and in both scenarios, the order of the paths is shuffled to avoid any ordering biases li2025generationjudgmentopportunitieschallenges .
+Selection Step.
+We construct a prompt for selecting the 10 most relevant paths with respect to the question-answer, and available optional answer pairs.
+Selection is intended to be done without any particular ordering, see Listing 8 in Appendix F.
+The set of paths for each question is processed in two passes, and in both scenarios, the order of the paths is shuffled to avoid any ordering biases (Li et al., 2025).
 From the two passes, we consider only the overlapping paths from the two candidate sets as the final collection of paths.
 During the selection phase, LLM-generated outputs are validated by checking for exact matches against the set of paths corresponding to the question.
-Any generated paths that do not match exactly are discarded to mitigate the risk of syntactic errors or hallucinations introduced by the LLM-as-a-judge.
+Any generated paths that do not have an exact match are discarded to mitigate the risk of syntactic errors or hallucinations introduced by the LLM-as-a-judge as a method of quality control.
 This process is repeated up to three times or until a total of 10 valid paths are obtained.
 If, after three attempts, fewer than 10 valid paths are selected, the remaining slots are filled by randomly sampling from the original KG path pool for the corresponding question.
-For questions that already contain 10 or fewer paths, the selection procedure is bypassed entirely.
-Scoring Step.Once a set of candidate paths is established, we construct another prompt for rating their relevance with respect to the given question and answer, and we process each path individually, see Appendix E Listing 9 for the instructions.
+The selection step is bypassed for questions that have 10 or fewer candidate paths.
+Scoring Step.
+Once a set of candidate paths is established, we construct another prompt for rating their relevance with respect to the given question and answer, and we process each path individually, see Appendix F Listing 9 for the instructions.
 Scoring is done by determining the quality score on a scale of 1-5, where 1 indicates a path which is completely unrelated to the question and answer, and 5 indicates an explicit answer to the question.
@@
 Table 2: Overview of Nllb200-3.3bn inference hyperparameters
-For enabling multilinguality for MultiHal, we employ the Nllb-200 3.3bn costa2022no model and focus on its five well-performing European languages, namely German, Italian, French, Portuguese and Spanish.
+For enabling multilinguality for MultiHal, we employ the Nllb-200 3.3bn (Costa-Jussà et al., 2022) model and focus on its five well-performing European languages, namely German, Italian, French, Portuguese and Spanish.
 Our generation hyperparameters are specified in Table 2.
@@
 3 Experimental Setup
-The baseline experiments are set up using a simple prompt-based knowledge injection method.
-The prompt $\mathcal{P}$ is formatted as $\mathcal{P}=(\mathcal{K},\mathcal{Q})$, where $\mathcal{K}$ is knowledge in the form of a KG path and $\mathcal{Q}$ is the question of the data point, see Appendix F Listing 10 for KG-RAG and Listing 11 for vanilla QA for used prompts.
-We conduct experiments with and without knowledge $\mathcal{K}$ (vanilla QA and KG-RAG respectively) to measure the effectiveness of the factual information contained in the KG paths.
-We measure the semantic similarity between ground-truths and model predictions using Multilingual-MiniLM-L12-v2666https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 wang2024multilingual , the choice of the sentence embedding model is based on results in the MMTE benchmark mmte_benchmark , its multilingual capabilities, and comparatively small parameter count.
+The baseline experiments are set up using a prompt-based knowledge injection method.
+The prompt $\mathcal{P}$ is formatted as $\mathcal{P}=(\mathcal{K},\mathcal{Q})$, where $\mathcal{K}$ is knowledge in the form of a KG path and $\mathcal{Q}$ is the question of the data point, see Appendix G Listing 10 for KG-RAG and Listing 11 for vanilla QA for used prompts.
+We conduct experiments with and without knowledge $\mathcal{K}$ (KG-RAG and vanilla QA respectively) to measure the effectiveness of the factual information contained in the KG paths.
+We measure the semantic similarity between ground-truths and model predictions using Multilingual-MiniLM-L12-v2666https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 (Wang et al., 2024), the choice of the sentence embedding model is based on results in the MMTE benchmark (Enevoldsen et al., 2025), its multilingual capabilities, and comparatively small parameter count.
 Semantic similarity is computed by mean-pooling the last hidden states per token for each sentence, applying L2 normalization and computing the dot-product between the ground-truth and LLM prediction representations.
 For experimental conditions, we use Gemini 2.0 Flash, GPT-4o Mini, and Llama 3.3 70bn instruct models.
+Additionally we triangulate our semantic similarity results with NLI and hallucination detection.
+Inspired from Bang et al.
+(2025) we run hallucination detection on the English subsplit of MultiHal using HHEM-2.1 (Bao et al., 2024), and run NLI evaluation similar to Sansford et al.
+(2024); Zhang et al.
+(2024) based on a DeBERTa multilingual NLI model  [^7] https://huggingface.co/MoritzLaurer/mDeBERTa-v3-base-xnli-multilingual-nli-2mil7 namely due to model’s language coverage and its performance.
+The NLI and HHEM-2.1 models expect test pairs as hypothesis and premise.
+We form the premise as a prompt containing a question and ground-truth $\mathcal{P}=\{\mathcal{Q};\mathcal{G}\}$ and hypothesis is model response $\mathcal{A}$.
 Additionally, we compute the Spearman correlation between the semantic similarity score and the quality score of each KG path, aiming to quantify the reliability of the quality score (see Section 2.3) determined by LLM-as-a-judge.
@@
 Considering that our methodology primarily relies on an LLM-judge for filtering and rating the quality of KG paths, we conduct a preliminary baseline test for the English subset to observe the performance of LLM judges.
-We conduct the experiment using a subset of MultiHal, see Figure 2 for the data distribution.
+We conduct the experiment using a proportionally sampled subset of MultiHal, see Figure 2 for the data distribution.
 The goal of this is to gain insights on the expected quality of the KG paths.
@@
 Table 3: Overview of preliminary baseline results for KG-RAG QA task with Gemini 2.0 Flash as answer generator. Results are based on dataset subsplits in Figure 2.
+Metric Value
+False Positives 11%
+False Negatives 2.78%
+IAA Cohen-Kappa 0.62
+Table 4: GPT 4o-Mini overview of false positives, false negatives and IAA - all computed with respect to human judgement.
+Before computation all path scores are binarized as low (ratings 1, 2, 3) and high (ratings 4, 5) quality.
+Results are based on dataset subsplits in Figure 2.
 From the results in Table 3 we see that paths judged by GPT 4o-Mini have a higher correlation with the semantic similarity.
-Assuming that low quality paths (rated 1-3) would impair the LLM output quality and high quality paths (4-5) would improve it, we chose to run the full baseline experiments with GPT 4o-Mini.
-For a more in-depth breakdown of GPT 4o-Mini performance per dataset and domain, please refer to Appendix G.
+Given that low quality paths (rated 1-3) impair the LLM output quality and high quality paths (4-5) improve it (see Appendix M), we chose to run the full baseline experiments with GPT 4o-Mini.
+For a more in-depth breakdown of GPT 4o-Mini performance per dataset and domain, please refer to Appendix H.
+Table 4 showcases false positives and false negatives for GPT 4o-Mini as well as interannotator agreement with respect to human judgment as ground truth.
+IAA score between human and GPT 4o-Mini annotations, including a focus on misclassifications by GPT-4o which gives an indication of noise presence which is comparable to datasets and can be used for reproducibility.
 4.2 Baseline Experiments
-We report our baseline experiment results in Figure 3.
-The results show the mean semantic similarity along with standard deviation across the data points for a given condition.
-For a full numerical overview, please refer to Appendix H Table 8.
-The results showcase a consistent improvement of KG-RAG setting over QA, indicating that the mined KG paths are meaningful for a model to generate a more semantically correct result than without.
-Given that Figure 3 showcases mean score improvements in the range of 0.12 - 0.36 points with comparable standard deviations, we assume the QA and KG-RAG results as seperate distributions, therefore we do not compute formal statistical significance tests.
-Table 4 presents more fine-grained results per domain and dataset for each test condition, with mean performance over all languages.
+We report our baseline experiment results in Figure 3 for semantic similarity, Table 6 for aggregated NLI scores over languages and Table 6 for hallucination detection on the English subsplit, see Appendix O for a fine-grained overview of semantic similarity over domains and Appendix N for NLI results expanded over each language.
+The results showcase a consistent improvement of KG-RAG setting over QA for all evaluation metrics, indicating that the mined KG paths are meaningful for a model to generate a higher quality output.
+The result distributions in Figure 3 have statistically significant differences across all languages and all models between the QA and KG-RAG conditions, see Appendix L for more details.
+For a full numerical overview of Figure 3, see Appendix I.
 Figure 3: Overview of the baseline experiment results showing mean semantic scores and standard deviation (as error bars) for QA and KG-RAG conditions over the whole MultiHal benchmark, seperated by language.
-Gemini 2.0 Flash Llama 3.3 70b instruct GPT 4o Mini
-Domain KG-Paths $\mathcal{Q}$ KG-RAG QA delta KG-RAG QA delta KG-RAG QA delta
-tqa_gen (misconceptions) 19 7 0.611 0.887 -0.276 0.754 0.803 -0.049 0.871 0.82 0.051
-tqa_gen (conspiracies) 29 6 0.588 0.854 -0.266 0.767 0.8 -0.033 0.844 0.833 0.011
-tqa_gen (paranormal) 4 2 0.495 0.724 -0.229 0.657 0.524 0.133 0.788 0.567 0.221
-tqa_gen (fiction) 7 5 0.45 0.545 -0.095 0.487 0.504 -0.017 0.569 0.511 0.058
-tqa_gen (indexical error: identity) 2 1 0.758 0.665 0.093 0.775 0.799 -0.024 0.956 0.97 -0.014
-tqa_gen (indexical error: time) 6 2 0.178 0.298 -0.12 0.281 0.334 -0.053 0.273 0.352 -0.079
-tqa_gen (indexical error: location) 1 1 0.053 0.294 -0.241 0.049 0.159 -0.11 0.065 0.195 -0.13
-tqa_gen (distraction) 10 5 0.389 0.191 0.198 0.421 0.343 0.078 0.446 0.373 0.073
-tqa_gen (advertising) 2 2 0.422 0.58 -0.158 0.576 0.671 -0.095 0.745 0.736 0.009
-tqa_gen (religion) 2 1 0.199 0.572 -0.373 0.426 0.68 -0.254 0.373 0.639 -0.266
-tqa_gen (stereotypes) 1 1 0.653 0.611 0.042 0.852 0.75 0.102 0.903 0.761 0.142
-tqa_gen (economics) 3 3 0.595 0.733 -0.138 0.72 0.808 -0.088 0.915 0.882 0.033
-tqa_gen (politics) 9 4 0.735 0.8 -0.065 0.844 0.822 0.022 0.835 0.828 0.007
-tqa_gen (law) 1 1 0.212 0.551 -0.339 0.534 0.643 -0.109 0.508 0.689 -0.181
-tqa_gen (language) 2 1 0.406 0.682 -0.276 0.588 0.683 -0.095 0.674 0.619 0.055
-tqa_gen (confusion: people) 15 7 0.639 0.312 0.327 0.598 0.294 0.304 0.575 0.251 0.324
-tqa_gen (confusion: places) 34 10 0.777 0.69 0.087 0.761 0.519 0.242 0.711 0.506 0.205
-tqa_gen (sociology) 10 3 0.624 0.82 -0.196 0.728 0.726 0.002 0.851 0.798 0.053
-tqa_gen (confusion: other) 1 1 0.819 0.291 0.528 0.752 0.459 0.293 0.568 0.575 -0.007
-tqa_gen (misinformation) 1 1 0.654 0.509 0.145 0.779 0.309 0.47 0.894 0.453 0.441
-tqa_gen (statistics) 1 1 0.329 0.696 -0.367 0.713 0.685 0.028 0.856 0.768 0.088
-tqa_gen (health) 2 2 0.348 0.589 -0.241 0.518 0.549 -0.031 0.555 0.612 -0.057
-tqa_gen (history) 20 5 0.555 0.705 -0.15 0.668 0.688 -0.02 0.705 0.686 0.019
-tqa_gen (nutrition) 1 1 0.606 0.761 -0.155 0.709 0.674 0.035 0.735 0.733 0.002
-tqa_gen (mandela effect) 6 3 0.526 0.566 -0.04 0.723 0.694 0.029 0.644 0.711 -0.067
-tqa_gen (logical falsehood) 4 1 0.256 0.796 -0.54 0.865 0.848 0.017 0.995 0.944 0.051
-defan (entertainment) 2803 556 0.868 0.547 0.321 0.802 0.469 0.333 0.74 0.499 0.241
-defan (nobleprize) 2718 557 0.764 0.769 -0.005 0.765 0.743 0.022 0.763 0.651 0.112
-defan (sports) 404 75 0.645 0.567 0.078 0.54 0.491 0.049 0.479 0.49 -0.011
-defan (worldorg) 875 118 0.689 0.304 0.385 0.38 0.277 0.103 0.273 0.259 0.014
-defan (qsranking) 3169 669 0.71 0.298 0.412 0.389 0.19 0.199 0.34 0.226 0.114
-felm (wk) 73 17 0.63 0.794 -0.164 0.742 0.789 -0.047 0.853 0.826 0.027
-halubench (general) 19 8 0.646 0.372 0.274 0.517 0.253 0.264 0.457 0.278 0.179
-halubench (pubmed) 586 182 0.167 0.645 -0.478 0.553 0.65 -0.097 0.689 0.713 -0.024
-halubench (finance) 6 3 0.203 0.564 -0.361 0.544 0.627 -0.083 0.643 0.653 -0.01
-halubench (covid) 15 7 0.696 0.681 0.015 0.692 0.577 0.115 0.716 0.632 0.084
-halueval (qa) 11398 3420 0.776 0.539 0.237 0.617 0.409 0.208 0.517 0.386 0.131
-shroom2024 (N/A) 346 160 0.454 0.371 0.083 0.511 0.447 0.064 0.469 0.471 -0.002
-simpleqa (geography) 433 153 0.72 0.391 0.329 0.54 0.3 0.24 0.428 0.259 0.169
-simpleqa (politics) 705 238 0.702 0.357 0.345 0.588 0.283 0.305 0.415 0.247 0.168
-simpleqa (other) 280 121 0.671 0.34 0.331 0.541 0.243 0.298 0.405 0.215 0.19
-simpleqa (science and technology) 848 304 0.709 0.316 0.393 0.607 0.246 0.361 0.431 0.2 0.231
-simpleqa (tv shows) 70 31 0.676 0.302 0.374 0.532 0.242 0.29 0.38 0.204 0.176
-simpleqa (music) 201 79 0.652 0.34 0.312 0.563 0.266 0.297 0.417 0.24 0.177
-simpleqa (art) 459 181 0.64 0.369 0.271 0.565 0.275 0.29 0.415 0.244 0.171
-simpleqa (sports) 169 87 0.622 0.31 0.312 0.538 0.243 0.295 0.381 0.199 0.182
-simpleqa (history) 109 46 0.621 0.384 0.237 0.567 0.27 0.297 0.399 0.224 0.175
-simpleqa (video games) 26 6 0.677 0.482 0.195 0.722 0.472 0.25 0.545 0.443 0.102
-Table 4: Breakdown of results per domains.
-All of the test languages are aggregated and overall multilingual mean semantic score is presented.
-Improvements are marked as bold.
-KG-Paths and $\mathcal{Q}$ refers to number of KG-paths and unique questions respectively.
+Model Task Ent Neut Contr
+Gemini-2.0 Flash KG-RAG 68% 26% 6%
+QA 52% 24% 24%
+Llama-3.3-70b instr KG-RAG 71% 21% 8%
+QA 45% 31% 23%
+GPT-4o Mini KG-RAG 76% 16% 8%
+QA 40% 32% 28%
+Table 5: Aggregated NLI results over all MultiHal languages. Ent is entailment, Neut is neutral, Contr is contradicting
+Model Task C H
+Gemini-2.0 Flash KG-RAG 87% 13%
+QA 58% 42%
+Llama-3.3-70b instr KG-RAG 88% 12%
+QA 55% 45%
+GPT-4o Mini KG-RAG 89% 11%
+QA 47% 53%
+Table 6: Result overview with hallucination detection using HHEM-2.1. C (consistent), H (hallucinated)
 5 Discussion
-Overall, the results from Figure 3 depict a consistent improvement for all tested LLMs in QA and KG-RAG settings; for detailed results refer to Table 8 in Appendix H.
+Overall, the results from Figure 3 and Tables 6, 6 depict a consistent improvement in a KG-RAG for all tested LLMs over QA; for detailed results of semantic similarity over domains refer to Appendix O.
 Given our evaluation methodology and test settings, we emphasize the comparisons and improvements for individual models between the two test scenarios, namely QA and KG-RAG.
@@
 Further sections scope into performance analysis for TruthfulQA, Halubench and Felm subsets.
-We supply some example problematic data points in Appendix I.
+We supply some example problematic data points in Appendix J.
+Table 4 reports approximately 11% false positives which indicates levels of noise although this is still comparable and exceeds other QA datasets where upon analysis the noise ranges in 20-30% (Iqbal et al., 2024).
 Temporal, Leading, Suggestive and Reasoning Questions.
@@
 Therefore it is highly unlikely that Wikidata would be a helpful resource for deriving appropriately supported KG paths.
-Our evaluation pipeline could benefit from integration of a reasoning methodology similar as per Generate-on-Graph genOnGraph or Think-on-Graph sun2023think .
-Refer to Appendix I Listings 15 and 13 for explicit examples.
+Our evaluation pipeline could benefit from integration of a reasoning methodology similar as per Generate-on-Graph (Xu et al., 2024) or Think-on-Graph (Sun et al., 2023).
+Refer to Appendix J Listings 15 and 13 for explicit examples.
 Domains.
 Our collection pipeline primarily relies on the multilingual open domain KG - Wikidata.
-For domains such as HaluBench-Pubmed, TruthfulQA-Health and HaluBench-Covid, performance can be improved by utilizing medical domain knowledge graphs, for example PubMed xu2020building or PrimeKG chandak2023building .
+For domains such as HaluBench-Pubmed, TruthfulQA-Health and HaluBench-Covid, performance can be improved by utilizing medical domain knowledge graphs, for example PubMed (Xu et al., 2020) or PrimeKG (Chandak et al., 2023).
 We outline that the modularity of our pipeline allows for easy substitute of the KG endpoint.
@@
 We also note that our sentence embedding evaluation may not always accurately capture the semantics with respect to the question.
-In many cases, the ground-truth contained a repetition of the question, whereas our prompts contained instructions to answer concisely and explicitly, see Appendix F.
+In many cases, the ground-truth contained a repetition of the question, whereas our prompts contained instructions to answer concisely and explicitly, see Appendix G.
 Consequentially, some data points were evaluated with a relatively low semantic score even though the model responses directly, or with minimal deviations, answered the question.
 We note that TruthfulQA and Felm have been particularly affected by this as their ground-truth answers contain repetitions of the text but our model responses are more focused on single, explicit entities without linguistic overheads.
-Refer to Appendix I Listings 12 and 14
+Refer to Appendix J Listings 12 and 14.
+Therefore we triangulate our semantic similarity results with NLI and hallucination detection for increased confidence in our benchmark quality.
 6 Related Work: Use of KGs in Datasets and Language Modeling
-Multiple surveys discuss KG usage in the context of LLMs, particularly outlining future work roadmaps and synergy russaKgSurvey ; pan2024unifying ; kau2024combining , discussing KGs in context of factuality, hallucination mitigation, multilinguality lavrinovics , and graph-retrieval augmented generation peng2024graph .
+Multiple surveys discuss KG usage in the context of LLMs, particularly outlining future work roadmaps and synergy (Pan et al., 2023; 2024; Kau et al., 2024), discussing KGs in context of factuality, hallucination mitigation, multilinguality (Lavrinovics et al., 2025), and graph-retrieval augmented generation (Peng et al., 2024).
 We identify these as useful starting points for researchers new to the topic.
 Language Modeling.
-Previous works make use of KG structures sansford2024grapheval ; rashad-etal-2024-factalign as part of their hallucination detection methodology by extracting graph structures from a given piece of text passage.
-Another approach involves reasoning over KGs sun2023think or generating SPARQL queries from natural text srivastava2024mst5 .
-FactKG kim-etal-2023-factkg and Fleek fatahi-bayat-etal-2023-fleek propose methodologies using KGs to aid fact-checking.
+Sansford et al.
+(2024); Rashad et al.
+(2024) make use of KG structures as part of their hallucination detection methodology by extracting graph structures from a given piece of text passage.
+Sun et al.
+(2023) approach involves reasoning over KGs and Srivastava et al.
+(2024) generates SPARQL queries from natural text.
+FactKG (Kim et al., 2023) and Fleek (Fatahi Bayat et al., 2023) propose methodologies using KGs to aid fact-checking.
 All the aforementioned language modeling approaches present KG information as in-context knowledge.
 However, in-context knowledge has limitations — particularly when there are conflicts between LLM’s internal knowledge and the provided context, or when there is limited transparency into how the model integrates and utilizes the external knowledge.
-An alternative approach is to encode the information as part of the model’s weights using adapter networks pfeiffer2023modular ; tian-etal-2024-kg ; ribeiro-etal-2022-factgraph .
+An alternative approach is to encode the information as part of the model’s weights using adapter networks (Pfeiffer et al., 2023; Tian et al., 2024; Ribeiro et al., 2022).
 Factually Oriented and KG-QA Datasets.
 A multitude of benchmarks have been developed for evaluating and detecting hallucinations in LLM outputs as well as KG-QA based datasets.
-Benchmarks such as Shroom2025 shroom2025 , Felm zhao2023felm , TruthfulQA lin2021truthfulqa , simpleQA , HaluBench halubench , HaluEval li2023halueval , DefAn rahman2024defan and SimpleQA simpleQA are intended for factuality evaluation of LLMs, consisting of different types of questions such as reasoning, information retrieval and they vary in domains.
+Benchmarks such as Shroom2025 (Vázquez et al., 2025), Felm (Zhao et al., 2023), TruthfulQA (Lin et al., 2021), (Wei et al., 2024), HaluBench (Ravi et al., 2024), HaluEval (Li et al., 2023), DefAn (Rahman et al., 2024) and SimpleQA (Wei et al., 2024) are intended for factuality evaluation of LLMs, consisting of different types of questions such as reasoning, information retrieval and they vary in domains.
 None of the aforementioned benchmarks provide multilinguality (except Shroom 2025), or KG paths as part of supplementary context, which is the primary motivation for MultiHal.
-Furthermore GRAF cruaciun2024graf is a legal domain KG-based benchmark for Romanian language, although is limited by lack of multilinguality.
-MintakaQA sen-etal-2022-mintaka and MKQA mkqa datasets offers multilingual coverage as well as annotations of Wikidata entities for questions-answers sen-etal-2022-mintaka or only answers mkqa , but not full KG paths.
+Furthermore GRAF (Crăciun et al., 2024) is a legal domain KG-based benchmark for Romanian language, although is limited by lack of multilinguality.
+MintakaQA (Sen et al., 2022) and MKQA (Longpre et al., 2021) datasets offers multilingual coverage as well as annotations of Wikidata entities for questions-answers (Sen et al., 2022) or only answers (Longpre et al., 2021), but not full KG paths.
 7 Conclusions
@@
 MultiHal is based around a multilingual question-answering task grounded with factual information; however ignoring use cases of multi-round dialogue and text summarization.
-Furthermore, our multilinguality can be considered limited in typological diversity ploeger-etal-2024-typological and bias towards European-centric languages.
-Multi-prompt evaluation has also been raised as an important component of LLM evaluation methodology whatSota ; maia2024efficient , which we do not cover, yet acknowledge as necessary, to test for LLM robustness and safety.
-Therefore, we propose expanding MultiHal’s downstream tasks and adding multiprompt evaluation and more diverse language coverage.
+Furthermore, our multilinguality can be considered limited in typological diversity (Ploeger et al., 2024).
+We do not include a multi-prompt evaluation (Mizrahi et al., 2024; Maia Polo et al., 2024) and leave it for future expansion of this benchmark.
 For evaluation of baseline experiments, we use three seperate models with no re-runs of random seeds.
 The evaluation of semantic similarity on a continuous scale makes the results hard to interpret across models, though still valid on a relative scale per model.
-Therefore, we propose future works of evaluating semantic similarity by framing it as a classification problem using a natural language inference sansford2024grapheval , or LLM-as-a-judge min-etal-2023-factscore .
+None of our evaluation metrics provide a fine-grained overview pinpointing exact hallucinatory text spans.
 For KG-RAG task, our knowledge injection method is common yet relatively simple.
 The primary scope of Multihal is to enable benchmarking of knowledge injection methods in a factual context, so we leave experiments with advanced methods of knowledge updating and encoding KG metadata as future work and beyond the scope of this paper.
+9 Reproducibility Statement
+We outline that majority of our dataset collection pipeline uses interpretable methods such as SPARQL queries or open-source frameworks such as Falcon 2.0.
+Our automated path quality evaluation is based on LLM judge methodology using GPT-4o-Mini which is closed source, although we report interannotator agreement and false positives with respect to human judgement.
+In case of reproducibility, these numbers can be used as a guideline for evaluating reproduced collection with open sourced models.
+All of our evaluation metrics use open sourced models.
+We also open-source our data collection codebase which can be easily modified by replacing closed-source components for open-source ones and we diligently conduct a multi-faceted analysis which can be used as reference.
+10 Ethics Statement
+To the best of our knowledge our work does not raise any ethical concerns with respect to ICLR ethics policy.
+We comply with the dataset licenses and terms of use.
+To the best of our knowledge, the datasets do not contain personally identifiable information or sensitive content that could raise ethical concerns.
+See Appendix B for our LLM usage statement.
 Appendix A MultiHal: Statistics and Schema
@@
 Table 8: MultiHal dataset schema.
-Appendix B Refusal Patterns
+Appendix B LLM Usage
+In writing of this paper we use LLMs as a writing aid for formatting, grammar correction and paraphrasing. All the LLM generated writing is human cross-checked and cross-validated.
+Appendix C Refusal Patterns
 List of refusal patterns used to filter data points with matching ground-truth answers.
 refusal_strings = ["I’m an AI", "I have no comment", "As an AI language model", "I am an", "I do not have", "I don’t have", "I am an artificial intelligence", "Nothing happens", "nothing in particular"]
-Appendix C SPARQL Queries
+Appendix D SPARQL Queries
 This section describes all the SPARQL queries used within the dataset gathering.
@@
 VALUES ?p4 {\$LIST_OF_TIMED_PROPERTIES}
-Listing 3: SPARQL query for 2-hop date retrieval subject-object. The \$OBJECT is a date string formatted as yyyy-mm-dd.
+Listing 3: SPARQL query for 2-hop date retrieval subject-object. The \$OBJECT is a date string formatted as yyyy-$m$-$d$.
 SELECT ?p1 ?o1 ?p2 ?o2 ?p3 ?o3 ?p4 ?o4 ?o99 WHERE {
@@
 Listing 5: SPARQL query for retrieving an entity label.
-Appendix D Numerical and Time-Based Properties
+Appendix E Numerical and Time-Based Properties
 time_properties = [’P569’, P570’, P571’, P574’, P575’, P576’, P577’, P580’, P582’, P585’, P606’, P619’, P620’, P621’, P622’, P729’, P730’, P746’, P813’, P1191’, P1249’, P1319’, P1326’, P1619’, P2285’, P2669’, P2913’, P3893’, P3999’, P5204’, P6949’, P7103’, P7104’, P7124’, P7125’, P7588’, P7589’, P8554’, P8555’, P8556’, P9052’, P9448’, P9667’, P10135’, P12044’, P12413’, P12506’, P12643’, P12686’, P12687’]
@@
 Listing 7: List of numerical properties.
-Appendix E LLM Judge Prompts
+Appendix F LLM Judge Prompts
 <instructions>
@@
 Listing 9: Prompt used for LLM-Judge KG path quality ratings.
-Appendix F Baseline Experiment Prompts
+Appendix G Baseline Experiment Prompts
 <instructions>
@@
 Listing 11: Prompt used for KG-RAG evaluation.
-Appendix G Overview of Preliminary Baseline Results per Domain and Dataset
+Appendix H Overview of Preliminary Baseline Results per Domain and Dataset
 Dataset Domain Num data points Mean Sem Score Mean Judged Score
@@
 Table 9: Breakdown of results of GPT-4o Mini from Table 3
-Appendix H Baseline Result Numerical Values
+Appendix I Baseline Result Numerical Values
+See Table 10 for numerical overview of our semantic similarity scores.
 Eng Deu Fra Ita Spa Por
@@
 QA performs vanilla question answering whereas KG-RAG provides mined KG paths as part of the input prompt for knowledge injection.
-Appendix I Example Problematic Data Points from Evaluation
+Appendix J Example Problematic Data Points from Evaluation
 We present a set of questions that serve as an example of the f
@@
 Listing 15: Questions of Suggestive, Leading or with Implicit Dependencies qualities
-Appendix J CO2 Emission and Compute Resources Related to Experiments
+Appendix K CO2 Emission and Compute Resources Related to Experiments
 We present an overview of our computation times for each of the core processing steps in Table 11.
@@
 Total emissions are estimated to be 1.19 kgCO2eq of which 0 percents were directly offset.
-Estimations were conducted using the MachineLearning Impact calculator presented in [16].
+Estimations were conducted using the MachineLearning Impact calculator presented in (Lacoste et al., 2019).
+Appendix L Statistical Significance Tests
+We compute Shapiro-Wilk (SW) test for semantic score normality distribution and Cramer-von-Misses two-sample test for statistical significance between QA and KG-RAG distribution means of full MultiHal benchmark, see Table 12.
+Language Model SW p-Value (KG-RAG) SW p-Value (QA) CvM p-Value
+eng Gemini 2.0 1.53e-110 3.50e-76 2.54e-07
+eng Llama 3.3 70B 2.48e-106 3.08e-71 3.40e-07
+eng GPT 4o-Mini 1.02e-77 9.02e-71 1.47e-07
+deu Gemini 2.0 3.23e-90 3.05e-73 1.01e-07
+deu Llama 3.3 70B 1.13e-74 5.61e-69 1.07e-07
+deu GPT 4o-Mini 7.94e-59 3.22e-63 8.26e-08
+fra Gemini 2.0 4.58e-81 1.36e-69 1.39e-07
+fra Llama 3.3 70B 6.84e-67 8.94e-70 1.28e-07
+fra GPT 4o-Mini 5.01e-56 6.17e-68 6.68e-08
+ita Gemini 2.0 4.84e-99 8.57e-75 2.34e-07
+ita Llama 3.3 70B 3.09e-76 2.97e-65 1.24e-07
+ita GPT 4o-Mini 3.38e-59 2.51e-62 1.17e-07
+spa Gemini 2.0 2.02e-96 1.67e-72 1.38e-07
+spa Llama 3.3 70B 1.94e-59 7.03e-64 9.78e-08
+spa GPT 4o-Mini 5.34e-60 5.40e-63 7.37e-08
+por Gemini 2.0 1.07e-104 3.20e-71 2.91e-07
+por Llama 3.3 70B 3.65e-76 3.67e-69 1.61e-07
+por GPT 4o-Mini 4.49e-65 5.20e-64 9.70e-08
+Table 12: Shapiro-Wilk (SW) and Cramer-von Mises (CvM) by Model and Language for distribution of semantic scores.
+Appendix M Path Quality Ablations
+We compute further results on ablating the KG path quality levels from lowest quality (1) up to highest quality (5).
+Table 13 shows gradual shift towards high quality paths.
+The results are computed over the preliminary baseline distribution as per Figure 2.
+Quality Levels Mean Std N
+[1] 0.33 0.28 455
+[1, 2] 0.37 0.31 1623
+[1, 2, 3] 0.42 0.33 2046
+[1, 2, 3, 4] 0.47 0.35 2451
+[1, 2, 3, 4, 5] 0.51 0.36 2834
+[2, 3, 4, 5] 0.55 0.36 2379
+[3, 4, 5] 0.70 0.32 1211
+[4, 5] 0.77 0.28 788
+[5] 0.81 0.26 383
+Table 13: Overview of semantic similarity with respect to KG path quality.
+Appendix N NLI Test Results
+Table 14 showcases NLI results expanded over languages with finer precission decimals with an aggregated total mean for a general overview.
+Name Language Task Entailment Neutral Contradiction
+google-gemini-2.0-flash-001 eng KG-RAG 65.70% 30.05% 4.25%
+QA 56.13% 19.12% 24.76%
+deu KG-RAG 67.79% 25.76% 6.45%
+QA 49.96% 25.69% 24.35%
+fra KG-RAG 58.32% 33.63% 8.06%
+QA 44.95% 33.72% 21.33%
+ita KG-RAG 75.53% 18.41% 6.06%
+QA 53.79% 19.84% 26.37%
+spa KG-RAG 71.86% 20.98% 7.16%
+QA 54.05% 23.45% 22.50%
+por KG-RAG 66.05% 28.28% 5.67%
+QA 52.70% 23.55% 23.75%
+meta-llama-llama-3.3-70b-instruct eng KG-RAG 68.91% 24.45% 6.64%
+QA 48.50% 30.08% 21.42%
+deu KG-RAG 71.31% 20.94% 7.76%
+QA 44.33% 31.80% 23.88%
+fra KG-RAG 64.70% 26.19% 9.11%
+QA 42.08% 36.44% 21.47%
+ita KG-RAG 74.36% 17.96% 7.68%
+QA 45.63% 28.55% 25.82%
+spa KG-RAG 71.01% 21.15% 7.84%
+QA 43.61% 32.31% 24.08%
+por KG-RAG 75.12% 18.09% 6.80%
+QA 48.04% 29.12% 22.84%
+openai-gpt-4o-mini eng KG-RAG 81.74% 12.10% 6.15%
+QA 42.70% 26.33% 30.97%
+deu KG-RAG 72.99% 18.42% 8.59%
+QA 37.61% 35.84% 26.55%
+fra KG-RAG 65.49% 24.48% 10.03%
+QA 37.22% 39.77% 23.00%
+ita KG-RAG 78.50% 13.65% 7.85%
+QA 41.14% 28.42% 30.44%
+spa KG-RAG 75.53% 15.74% 8.73%
+QA 40.77% 32.31% 26.92%
+por KG-RAG 80.93% 12.38% 6.69%
+QA 41.44% 31.38% 27.18%
+Total KG-RAG 71.44% 21.26% 7.31%
+QA 45.81% 29.32% 24.87%
+Table 14: NLI results over MultiHal benchmark.
+Appendix O Fine-Grained Semantic Similarity Overview Across Domains
+Table 15 showcases a breakdown of semantic similarity across domains of MultiHal.
+Gemini 2.0 Flash Llama 3.3 70b instruct GPT 4o Mini
+Domain KG-Paths $\mathcal{Q}$ KG-RAG QA delta KG-RAG QA delta KG-RAG QA delta
+tqa_gen (misconceptions) 19 7 0.611 0.887 -0.276 0.754 0.803 -0.049 0.871 0.82 0.051
+tqa_gen (conspiracies) 29 6 0.588 0.854 -0.266 0.767 0.8 -0.033 0.844 0.833 0.011
+tqa_gen (paranormal) 4 2 0.495 0.724 -0.229 0.657 0.524 0.133 0.788 0.567 0.221
+tqa_gen (fiction) 7 5 0.45 0.545 -0.095 0.487 0.504 -0.017 0.569 0.511 0.058
+tqa_gen (indexical error: identity) 2 1 0.758 0.665 0.093 0.775 0.799 -0.024 0.956 0.97 -0.014
+tqa_gen (indexical error: time) 6 2 0.178 0.298 -0.12 0.281 0.334 -0.053 0.273 0.352 -0.079
+tqa_gen (indexical error: location) 1 1 0.053 0.294 -0.241 0.049 0.159 -0.11 0.065 0.195 -0.13
+tqa_gen (distraction) 10 5 0.389 0.191 0.198 0.421 0.343 0.078 0.446 0.373 0.073
+tqa_gen (advertising) 2 2 0.422 0.58 -0.158 0.576 0.671 -0.095 0.745 0.736 0.009
+tqa_gen (religion) 2 1 0.199 0.572 -0.373 0.426 0.68 -0.254 0.373 0.639 -0.266
+tqa_gen (stereotypes) 1 1 0.653 0.611 0.042 0.852 0.75 0.102 0.903 0.761 0.142
+tqa_gen (economics) 3 3 0.595 0.733 -0.138 0.72 0.808 -0.088 0.915 0.882 0.033
+tqa_gen (politics) 9 4 0.735 0.8 -0.065 0.844 0.822 0.022 0.835 0.828 0.007
+tqa_gen (law) 1 1 0.212 0.551 -0.339 0.534 0.643 -0.109 0.508 0.689 -0.181
+tqa_gen (language) 2 1 0.406 0.682 -0.276 0.588 0.683 -0.095 0.674 0.619 0.055
+tqa_gen (confusion: people) 15 7 0.639 0.312 0.327 0.598 0.294 0.304 0.575 0.251 0.324
+tqa_gen (confusion: places) 34 10 0.777 0.69 0.087 0.761 0.519 0.242 0.711 0.506 0.205
+tqa_gen (sociology) 10 3 0.624 0.82 -0.196 0.728 0.726 0.002 0.851 0.798 0.053
+tqa_gen (confusion: other) 1 1 0.819 0.291 0.528 0.752 0.459 0.293 0.568 0.575 -0.007
+tqa_gen (misinformation) 1 1 0.654 0.509 0.145 0.779 0.309 0.47 0.894 0.453 0.441
+tqa_gen (statistics) 1 1 0.329 0.696 -0.367 0.713 0.685 0.028 0.856 0.768 0.088
+tqa_gen (health) 2 2 0.348 0.589 -0.241 0.518 0.549 -0.031 0.555 0.612 -0.057
+tqa_gen (history) 20 5 0.555 0.705 -0.15 0.668 0.688 -0.02 0.705 0.686 0.019
+tqa_gen (nutrition) 1 1 0.606 0.761 -0.155 0.709 0.674 0.035 0.735 0.733 0.002
+tqa_gen (mandela effect) 6 3 0.526 0.566 -0.04 0.723 0.694 0.029 0.644 0.711 -0.067
+tqa_gen (logical falsehood) 4 1 0.256 0.796 -0.54 0.865 0.848 0.017 0.995 0.944 0.051
+defan (entertainment) 2803 556 0.868 0.547 0.321 0.802 0.469 0.333 0.74 0.499 0.241
+defan (nobleprize) 2718 557 0.764 0.769 -0.005 0.765 0.743 0.022 0.763 0.651 0.112
+defan (sports) 404 75 0.645 0.567 0.078 0.54 0.491 0.049 0.479 0.49 -0.011
+defan (worldorg) 875 118 0.689 0.304 0.385 0.38 0.277 0.103 0.273 0.259 0.014
+defan (qsranking) 3169 669 0.71 0.298 0.412 0.389 0.19 0.199 0.34 0.226 0.114
+felm (wk) 73 17 0.63 0.794 -0.164 0.742 0.789 -0.047 0.853 0.826 0.027
+halubench (general) 19 8 0.646 0.372 0.274 0.517 0.253 0.264 0.457 0.278 0.179
+halubench (pubmed) 586 182 0.167 0.645 -0.478 0.553 0.65 -0.097 0.689 0.713 -0.024
+halubench (finance) 6 3 0.203 0.564 -0.361 0.544 0.627 -0.083 0.643 0.653 -0.01
+halubench (covid) 15 7 0.696 0.681 0.015 0.692 0.577 0.115 0.716 0.632 0.084
+halueval (qa) 11398 3420 0.776 0.539 0.237 0.617 0.409 0.208 0.517 0.386 0.131
+shroom2024 (N/A) 346 160 0.454 0.371 0.083 0.511 0.447 0.064 0.469 0.471 -0.002
+simpleqa (geography) 433 153 0.72 0.391 0.329 0.54 0.3 0.24 0.428 0.259 0.169
+simpleqa (politics) 705 238 0.702 0.357 0.345 0.588 0.283 0.305 0.415 0.247 0.168
+simpleqa (other) 280 121 0.671 0.34 0.331 0.541 0.243 0.298 0.405 0.215 0.19
+simpleqa (science and technology) 848 304 0.709 0.316 0.393 0.607 0.246 0.361 0.431 0.2 0.231
+simpleqa (tv shows) 70 31 0.676 0.302 0.374 0.532 0.242 0.29 0.38 0.204 0.176
+simpleqa (music) 201 79 0.652 0.34 0.312 0.563 0.266 0.297 0.417 0.24 0.177
+simpleqa (art) 459 181 0.64 0.369 0.271 0.565 0.275 0.29 0.415 0.244 0.171
+simpleqa (sports) 169 87 0.622 0.31 0.312 0.538 0.243 0.295 0.381 0.199 0.182
+simpleqa (history) 109 46 0.621 0.384 0.237 0.567 0.27 0.297 0.399 0.224 0.175
+simpleqa (video games) 26 6 0.677 0.482 0.195 0.722 0.472 0.25 0.545 0.443 0.102
+Table 15: Breakdown of results per domains.
+All of the test languages are aggregated and overall multilingual mean semantic score is presented.
+Improvements are marked as bold.
+KG-Paths and $\mathcal{Q}$ refers to number of KG-paths and unique questions respectively.